### v1

# STaRK: Benchmarking LLM Retrieval on Textual and Relational Knowledge Bases

Abstract
Answering real-world user queries, such as product search, often requires accurate retrieval of information from semi-structured knowledge bases or databases that involve blend of unstructured (e.g., textual descriptions of products) and structured (e.g., entity relations of products) information. However, previous works have mostly studied textual and relational retrieval tasks as separate topics. To address the gap, we develop STaRK, a large-scale Semi-structure retrieval benchmark on Textual and Relational Knowledge Bases. We design a novel pipeline to synthesize natural and realistic user queries that integrate diverse relational information and complex textual properties, as well as their ground-truth answers. Moreover, we rigorously conduct human evaluation to validate the quality of our benchmark, which covers a variety of practical applications, including product recommendations, academic paper searches, and precision medicine inquiries. Our benchmark serves as a comprehensive testbed for evaluating the performance of retrieval systems, with an emphasis on retrieval approaches driven by large language models (LLMs). Our experiments suggest that the STaRK datasets present significant challenges to the current retrieval and LLM systems, indicating the demand for building more capable retrieval systems that can handle both textual and relational aspects.
Figure 1: STaRK features queries on semi-structured knowledge bases (SKBs) with textual and relational knowledge, with node entities as ground truth answers. It simulates user interactions with a SKB to evaluate LLM retrieval systems’ performance in providing accurate responses.
Example query Title of ground truth node(s)
STaRK-Amazon Looking for durable Dart World brand dart flights that <Amazon Standard Flights>
resist easy tearing. Any recommendations? <Dart World Broken Glass Flight> (12 more)
What are recommended scuba diving weights for experienced divers <Sea Pearls Vinyl Coated Lace Thru Weight>
that would fit well with my Gorilla PRO XL waterproof bag?
STaRK-MAG Search publications by Hao-Sheng Zeng on non-Markovian dynamics. <Distribution of non-Markovian intervals…>
<Comparison between non-Markovian…>
What are some nanofluid heat transfer research papers published by <A Numerical Study on Convection Around A
scholars from Philadelphia University? Suqare Cylinder using AL2O3-H2O Nanofluid>
STaRK-Prime Could you provide a list of investigational drugs that <(S)-3-phenyllactic Acid>,
interact with genes or proteins active in the epididymal region? <Anisomycin>, <Puromycin>
Search for diseases without known treatments and induce <Intrahepatic Cholestasis>
pruritus in pregnant women, potentially associated with Autoimmune.
Please find pathways involving the POLR3D gene within nucleoplasm. <RNA Polymerase III Chain Elongation>
Which gene or protein associated with lichen amyloidosis can <OSMR>, <IL31RA>
bind interleukin-31 to activate the PI3K/AKT and MAPK pathways?
Table 1: Example queries from STaRK, which involve semi-structured information, where the relational and textual aspects are highlighted.
1 Introduction
Natural-language queries are the primary forms in human society for acquiring information, involving diverse knowledge in the real world [15, 23, 19]. For example, users on e-commerce web search can express complex information needs by combining free-form elements or constraints, such as “Can you help me find a push-along tricycle from Radio Flyer that’s both fun and safe for my kid?”. Here, the underlying knowledge can be represented in semi-structured formats [30, 41, 33], referred to as semi-structured knowledge bases (SKBs), which integrate unstructured data, such as textual descriptions and natural language expressions (e.g., descriptions about a tricycle), with structured data, like entity interactions on knowledge graph or database (e.g., a tricycle with “belongs to” relation with brand Radio Flyer). This allows semi-structured knowledge bases to represent comprehensive knowledge in specific applications, making them indispensable in domains such as e-commerce [13], social media [27], and biomedicine [6, 16]. In this work, we focus on semi-structured knowledge bases (SKBs) with texts and relation information, which are two prevalent modalities on the existing knowledge bases.
In this context, as shown in Figure 1, information retrieval serves as a predominant step by identifying relevant data from the semi-structured knowledge bases, such as specific tricycles that match the request, which are then processed to generate accurate answers to user queries [35]. The retrieval accuracy on semi-structured knowledge bases is crucial for enhancing user experience, supporting informed decision-making, and preventing hallucination.
Limitations of existing works. However, previous works and benchmarks have primarily focused on either purely textual queries on unstructured knowledge [18, 12, 22, 25, 44] or structured SQL or knowledge graph queries [2, 48, 46, 37], which are inadequate to study the complexities of retrieval tasks involving semi-structured knowledge bases. Another line of works [10, 21, 46, 42, 43, 38] have explored the middle ground via multi-hop reasoning within single or across multiple documents. However, they are either limited in the span of knowledge or lack of diverse textual properties and explicit relational information. Recently, large language models (LLMs) have demonstrated significant potential on information retrieval tasks [52, 12, 26, 34]. Nevertheless, it remains an open question how effectively LLMs can be applied to the specific challenge of retrieval from SKBs. It is also important to note that existing works focus mainly on general knowledge, e.g., from wikipedia. In fact, the knowledge may also come from private sources, necessitating information retrieval systems adaptable to private semi-structured knowledge bases. Therefore, there is a gap in our understanding of how current LLM retrieval systems handle the complex interplay between textual and relational requirements in queries that optionally involve private knowledge.
Challenges. To address this gap, our goal is to develop a benchmark for LLM retrieval on SKBs. However, the lack of real queries on SKBs has become an obstacle. Although retrieval datasets with synthetic queries is a valuable supplement, accurately simulating user queries on SKBs is particularly difficult. This difficulty arises from the interdependence of textual and relational information, which can lead to challenges in precisely filtering through millions of candidates to construct the ground truth answers. Additionally, ensuring that queries are relevant to real-world applications and resembles real-world scenarios adds further complexity to the benchmarking process.
Our contribution: STaRK. As shown in Figure 1, we present a large-scale Semi-structure retrieval benchmark on Textual and Relational Knowledge Bases (STaRK). We propose a novel pipeline that simulates user queries and constructs precise ground truth answers using three SKBs built from extensive texts and millions of entity relations from public sources. We validate the quality of queries in our benchmark through detailed analysis and human evaluation, focusing on their naturalness, diversity, and practicality. With STaRK, we delve deeper into retrieval tasks on SKBs, evaluate the capability of current retrieval systems, and providing insights for future advancement. We point out the key features of STaRK as follows:
  • Natural-sounding queries on semi-structured knowledge: As illustrated in Table 1, the queries in our benchmark are crafted to incorporate rich relational information and complex textual properties. Additionally, these queries closely mirror the types of questions users would naturally ask in real-life scenarios, e.g., with flexible query formats and possibly with extra contexts.
  • Context-specific reasoning: The queries entail reasoning capabilities specific to the context. This includes the ability to infer customer interests, understand specialized field descriptions, and deduce relationships involving multiple subjects mentioned within the query. For example, the context “I had a dozen 2.5-inch Brybelly air hockey pucks, so I’m trying to find matching strikers.” entails the user’s interest in looking for complementary products. Such reasoning capabilities are crucial for accurately interpreting and responding to the nuanced requirements of each query.
  • Diverse domains: Our benchmark spans a wide range of content and domains, reflecting the diverse nature of real-world information needs. It covers areas on product recommendation, academic paper search, and precision medicine inquiry. Therefore, STaRK provides a comprehensive evaluation of retrieval systems across various contexts and knowledge domains.
Moreover, we perform extensive experiments on each benchmark dataset for the prevailing retrieval systems. We highlight research challenges especially on the capability to handle textual and relational requirements and the latency of retrieval systems on large-scale SKBs that involve million-scale entities or relations. Finally, we provide insights into future directions to build more capable retrieval systems. The benchmark datsets and code are available on https://github.com/snap-stanford/stark.
2 Related Work
We identify three primary categories on retrieval-oriented question answering (QA) datasets based on the knowledge sources they utilize: unstructured, structured, and semi-structured knowledge bases. We discuss the differences and limitations in the previous works that we seek to address.
Unstructured QA Datasets. The line of work explores retrieving answers from a single document [32] or multiple document sources [40, 45, 21, 42, 10]. For example, SQuAD [32] focuses on extracting answers from a single document, testing model’s ability to understand and retrieve information within a specific context. For multi-document retrieval, HotpotQA [45] demands reasoning over multiple documents and includes sentence-level supporting facts, enhancing the complexity of the retrieval task. TriviaQA [21] combines questions from trivia games with evidence documents, evaluating both retrieval and reasoning over multiple sources. Moreover, some works [29, 24] leverage generated results by search engines as ground truth answers or information sources for retrieval. However, unstructured QA datasets or knowledge bases often lack the depth of relational reasoning commonly required in answering complex user queries.
Structured QA Datasets. Within this category, the datasets challenge models to retrieve answer from structured knowledge bases, either knowledge graphs [3, 49, 39, 4, 5, 11, 14] or tabular data [51, 50]. For example, ComplexWebQuestions [39] challenge model’s abilities to decompose complex queries and interpret queries using entities and relationships defined in knowledge graphs. GraphQA [14] textualizes graphs by describing node attributes and edges with natural language, offering language model a way to handle KBQA. For datasets with tabular data, WikiSQL [51] and Spider [50] both address the task of converting natural language questions to SQL queries, with WikiSQL concentrating on single-table databases and Spider extending to complex multi-tables. While these datasets primarily focus on relational information, the lack of textual information limits questions within predefined relationships and entities, which constrains the breadth of available information, resulting in less comprehensive and nuanced responses compared to the extensive insights drawn from abundant textual data.
Semi-Structured QA Datasets. Another line of work focuses on semi-structured QA with tabular and textual data. WikiTableQuestions [31] emphasizes comprehension of table structure and textual content, reflecting the semi-structured nature of many real-world datasets. TabFact [7] and HybridQA [8] combine textual and tabular data, requiring models to validate statements or answer questions using both data types. TabMCQ [20] introduces multiple-choice questions based on tables, adding complexity in understanding and reasoning over semi-structured data. However, different from this work, these existing datasets mainly focus on tables as structured sources, which may not incorporate rich relational information naturally exists between entities. Additionally, previous attempts to combine text and tabular data through external web links or additional text columns often result in data structures that are hard to navigate. Therefore, there is an emerging need to benchmark retrieval tasks on textual and relational knowledge base, offering more insights beyond the scope of structured tabular sources.
3 Benchmarking Retrieval Tasks over Textual and Relational Knowledge
Table 2: Data statistics of our constructed semi-structured knowledge bases
#entity #relation avg. #entities #relations #tokens
types types degree
STaRK-Amazon 2 3 3.0 1,032,407 3,886,603 592,067,882
STaRK-MAG 4 4 10.6 1,872,968 19,919,698 212,602,571
STaRK-Prime 10 18 62.6 129,375 8,100,498 31,844,769
Figure 2: Demonstration of Semi-structured Knowledge Bases, where each knowledge base combines both textual and relational information in a complex way, making the retrieval tasks challenging.
We develop three large-scale retrieval datasets on three semi-structure knowledge bases (SKBs). In Section 3.1, we introduce the schema and scale of the SKBs, which integrate entities with rich relational and textual information. In Section 3.2, we introduce the retrieval datasets and the key characteristics of the queries. Then, we highlight our novel dataset construction pipeline in Section 3.3, which simulates user queries involving textual and relational knowledge and automatically generates the ground truth answers in the retrieval datasets. Finally, we conduct data distribution analysis and human evaluation on the queries in Section 3.4.
3.1 Semi-structured Knowledge Bases
As shown in Figure 2, we construct three SKBs from the relational structure between entities and the textual information associated with a subset of the entities. We present the statistics of the relational structure in Table 2 and introduce each SKB as follows:
Amazon Semi-structured Knowledge Base. This knowledge base is derived from the Sports and Outdoors category from Amazon Product Reviews [13] and Amazon Question and Answer Data [28]. The knowledge base features two entity types: product and brand, and three types of relational information: also_bought, also_viewed between product entities, and has_brand relation between product and brand entities. In total, it comprises around 1.03M entities (product entities: 0.96M, brand entities: 0.07M) and 3.9M relations (also_bought: 1.7M, also_viewed: 1.3M, has_a_brand: 0.9M). We obtain the textual information by combining the meta data from Amazon Product Reviews with the customer Q&A records from Amazon Question and Answer Data. This provides a rich amount of textual data, including product titles, descriptions, prices, customer reviews, and related customer Q&A for each product. For brand entities, we extract brand titles as the textual attribute. Especially, Amazon SKB features an extensive textual data, which is largely contributed from customer reviews and Q&A.
MAG Semi-structured Knowledge Base. This knowledge base is constructed from obgn-MAG [17], obgn-papers100M [17], and Microsoft Academic Graph (version 2019-03-22) [41, 36]. The knowledge base contains around 1.9M entities under four entity types (author: 1.1M, paper: 0.7M, institution: 9K, field_of_study: 0.06M) and 20M relations under four relation types (author_writes_paper: 7.9M, paper_has_field_of_study: 7.2M, paper_cites_paper: 4.9M, author_affiliated_with_institution: 1.0M). We construct the knowledge base by selecting papers1 shared between obgn-MAG and obgn-papers100M. The relational structure comes from the extracted subgraph in obgn-mag based on the selected papers. The textual information such as titles and abstracts is sourced from obgn-papers100M. Additionally, we augment the textual data by integrating details from the Microsoft Academic Graph database, providing extra information like paper venue, author and institution names. This SKB demonstrates a large number of entities and relations associate with paper nodes, especially on citation and authorship relation types.
Prime Semi-structured Knowledge Base. We leverage the existing knowledge graph PrimeKG [6] which contains ten entity types including disease, drug, gene/protein, and eighteen relation types, such as associated_with, synergistic_interaction, indication, expression_present. The entity count in our knowledge base is approximately 129K, with around 8M relations. The details on entity numbers and relation tuples are available in Appendix A.2. Compared to the Amazon and MAG SKBs, Prime SKB is denser and features a greater variety of relation types. While PrimeKG already provides text information on disease and drug entities, including drug mechanisms and disease descriptions, we additionally integrate the textual details from multiple databases for gene/protein and pathway entities such as genomic position, gene activity summary and pathway orthologous event.
We include the comprehensive public sources we used to construct the SKBs in Table 8 of Appendix A.
footnotetext: 1 We filter out non-English language papers as we only considers single-lingual queries.
3.2 Retrieval Tasks on Semi-structured Knowledge Bases
Table 3: Dataset statistics on STaRK.
#queries #queries w/ avg. #answers train / val / test
multiple answers
STaRK-Amazon 9,100 7,082 17.99 0.65 / 0.17 / 0.18
STaRK-MAG 13,323 6,872 2.78 0.60 / 0.20 / 0.20
STaRK-Prime 11,204 4,188 2.56 0.55 / 0.20 / 0.25
We develop three novel retrieval datasets. Specifically, the task is to retrieve node entities over the SKBs given any input queries. To highlight, the queries feature the interplay and fusion between relational and textual knowledge. Moreover, to elevate their applicability in practical scenarios, these queries are designed to mimic real-life query patterns in terms of the natural-sounding property and flexible formats. Each query may have single answer or multiple answers, formulated as a subset of entities from the knowledge base, which requires the model to accurately identify the ground truth entities out of potentially millions in response to each query. In Table 1, we demonstrate example queries from these benchmarks. We detail our retrieval benchmarks’ scale in Table 3 and introduce the specifics of each retrieval dataset as follows:
STaRK-Amazon. This dataset comprises 9,100 customer-focused queries aimed at product searches, with a notable 68% of these queries yielding multiple answers. The dataset prioritizes customer-oriented criteria, highlighting textual elements such as product quality, functionality, and style. Additionally, it accentuates relational aspects, including brand and product connections (e.g., complementary or substitute items). The queries are framed in detailed, conversation-like formats, enriching the context and enhancing the dataset’s relevance to real-world scenarios.
STaRK-MAG. This dataset comprises 13,323 paper queries, with roughly half yielding multiple answers. Beyond the single-hop relational requirements, STaRK-MAG also emphasizes the fusion between the textual requirements with multi-hop queries. For example, “Are there any papers from King’s College London” highlights the metapath (institution $\rightarrow$ author $\rightarrow$ paper) on the relational structure. We designed three single-hop and four multi-hop relational query templates, cf. Appendix A.1. The textual aspects of the queries diversifies, focusing on different elements such as the paper’s topic, methodology, or the advancements it introduces, sourced mainly from the abstracts.
STaRK-Prime. This dataset, comprising 11,204 queries across all ten entity types, incorporates single-hop and multi-hop relational information similar to STaRK-MAG. We developed 28 multi-hop query templates, detailed in Appendix A.2, to cover various relation types and ensure their practical relevance. For instance, the query “What is the drug that targets the genes or proteins expressed in <anatomy>?” serves applications in precision medicine and pharmacogenomics, aiding researchers and healthcare professionals in identifying drugs that act on genes or proteins associated with specific anatomical areas and enabling more targeted treatments. For entities like drug, disease, gene/protein, and pathway, the queries are a hybrid of relational requirements and textual requirements drawn from documents. A notable feature is the deliberate simulation of three distinct roles – medical scientist, doctor, and patient – for certain queries related to drugs and diseases. This is intended to diversify the language used across various user types and to evaluate the robustness of the retrieval system under varied linguistic scenarios. For entities such as effect/phenotype, the queries rely solely on relational data due to limited textual information in the knowledge base.
We divide each dataset into training, validation, and testing subsets, with the ratios detailed in Table 3. For the STaRK-Amazon dataset, we randomly allocate a subset of queries with 20 or fewer answers to the validation and testing sets, while the remaining queries are assigned to the training set. For the STaRK-MAG and STaRK-Prime, we use random splits to divide the data.
3.3 Benchmark Construction
Figure 3: The process of constructing semi-structured retrieval datasets involves four main steps: 1) Sample Relational Requirement: Based on relational templates, sample a relational requirement on a SKB. 2) Extract Textual Properties: From a node that meets the relational requirement, extract relevant textual properties. 3) Combine Information: Merge the relational information and textual properties to form a natural-sounding query. 4) Construct Ground Truth Nodes: Check if nodes satisfy the textual properties using multiple language models to establish ground truth nodes.
Here we detail the novel pipeline to synthesize the retrieval datasets present in Section 3.2. The core idea is to entangle relational information and textual properties, while constructing ground truth answers in an accurate and efficient way. As shown in Figure 3, the construction of the retrieval datasets generally involves four steps, and the specific processes vary depending on the characteristics of each dataset. These steps are as follows:
  • Sample Relational Requirements: Initially, we sample a relation template, such as “(a product) belongs to <brand>” and instantiate it into a relational requirement, e.g., “belongs to Radio Flyer” shown in the first step of Figure 3. Each relational requirement yields a set of candidate entities that meet the requirement, i.e., products belonging to Radio Flyer. Therefore, we obtain the relational requirements and the corresponding candidates. The relation templates for the three SKBs are available in Appendix A, including 28 relation templates for the Prime SKB. Each relation template in the SKBs serves a distinct and practical purpose. For instance, the template “What is the drug that targets the genes or proteins which are expressed in <anatomy>?” is particularly valuable for precision medicine inquiries. It aids in pinpointing medications aimed at specific genetic or protein targets within defined anatomical areas, thereby contributing to the development of targeted treatments.
  • Extracting Textual Properties: Subsequently, we select one of the candidate nodes that satisfies the relational requirement, which is referred to as the gold answer. We extract textual properties that align with the interests of specific roles from the textual information of the gold answer. The roles could be customers, researchers, or medical scientists, differing across the SKBs. For example, in the second step of Figure 3, we extract the phrases of functionality and user experience from a Radio Flyer product, which come from different sources in the product document. We utilize GPT-3.5-turbo-16k for the STaRK-Amazon and Claude-v2 for the other two datasets to conduct this step, where the prompts used are included in Appendix B.
  • Combining Textual and Relational Information: With the textual and relational requirements, we simulate queries by fusing them using LLMs. We conduct two-stage fusion using two different LLMs, i.e., Claude-v2 and GPT-4-Turbo, to avoid bias that might arise from relying on a single LLM and ensure a more diverse set of simulated queries. Specifically, this first-stage integration is guided by various criteria, such as ensuring a natural-sounding query, adhering to the style of ArXiv searches, or aligning with specific roles. Furthermore, the second stage instructs the LLM to enrich the context and rephrase the language, thereby posing a more demanding reasoning challenge in comprehending the requirements of the query. In our example, by combining relational information “belongs to Radio Flyer” and textual information “push-along tricycle” and “fun and safe for kids”, we arrive at the final query in the third step.
  • Filtering Additional Answers: In the final step, we assess whether each remaining candidate, excluding the gold answer used for extracting textual properties, meets the additional textual requirement. For the verification, we employ three Claude models. Only candidates that pass the verification across all models are included in the final ground truth answer set. Therefore, this stringent verification ensures the accuracy of the nodes in the ground truth answers. During this process, we also evaluate the accuracy of the gold nodes that pass the verification criteria. The accuracy rates STaRK-Amazon, STaRK-MAG, and STaRK-Primeare 86.6%, 98.9%, and 92.3%, respectively, demonstrating the effectiveness of our filtering approach in maintaining high-quality ground truth answers.
This dataset construction pipeline is automatic, efficient, and broadly applicable to the SKBs within our context. We include the complete prompts of the above steps in Appendix B.
3.4 Benchmark Data Distribution Analysis and Human Evaluation
Figure 4: Distribution of query and answer lengths on STaRK datasets.
Table 4: Query diversity measurement on STaRK.
Shannon Type-Token
Entropy Ratio
STaRK-Amazon 10.39 0.179
STaRK-MAG 10.25 0.180
STaRK-Prime 9.63 0.143
Reference1 10.44 0.261
Figure 5: Average relative composition of relational vs. textual information.
We study the data distribution to understand the characteristics of our benchmark datasets and conduct human evaluation on the benchmark quality. We study the data distribution in three dimensions:
  • Query and Answer Length. We analyze the query length distribution (measured in the number of words) and the number of ground truth answers for each query. The query length reflects the amount of context information provided by users, while the number of ground truth answers indicates the level of inclusiveness or ambiguity of queries within the specific SKB. As illustrated in Figure 4, all three datasets exhibit similar query length distributions, with most queries containing around 16 words. Queries with up to 50 words typically involve mentions of other entities, such as product or paper titles, or provide more detailed context, such as specific symptoms descriptions. Interestingly, the answer length distribution for STaRK-Amazon shows a more significant long-tail pattern, with approximately 22% of answers exceeding 30 entities, whereas the answer lengths for STaRK-Prime and STaRK-MAG are all within 20 entities. On average, as shown in Table 3, STaRK-Amazon presents an average answer length of 5.32, indicative of the e-commerce recommendation domain where general customer inquiries frequently result in diverse recommendations. While STaRK-Prime has the smallest average answer length, partially attributed to the smaller size of the Prime SKB compared to the other two SKBs.
  • Query Diversity. A diverse set of queries pose challenges for broader applicability to meet varying user demands. Specifically, we measure query diversity using Shannon Entropy, which quantifies the uncertainty in the word distribution across all queries, and the Type-Token Ratio (TTR), which calculates the proportion of unique words to the total number of words. Higher values of Shannon Entropy and TTR indicate greater lexical diversity, reflecting a wider range of topics and linguistic expressions in the query set. As shown in Table 5, we observe high values of Shannon Entropy on all of the datasets. For reference, we compute the metrics for the wikipedia page of Barack Obama1footnotetext: 1https://en.wikipedia.org/wiki/Barack_Obama. While TTR is more sensitive to the document length, the TTR values consistently demonstrate a steady presence of unique words relative to the large length in our query document.
  • Ratio of Relational vs. Textual Information. A key feature of our benchmark dataset is the composition of textual and relational information. Therefore, it is crucial to understand the proportionality between these two types of information. We calculate the ratio of the length of relational requirements to the length of textual requirements for each query and then average these ratios across each dataset. Intuitively, the ratio serves as an approximation of the relative distribution based on the length of requirements, and does not directly reflect the importance of each type of information in determining the final answers. As shown in Figure 5, the ratios vary across different datasets, highlighting a differing emphasis on textual versus relational information. This distribution variation presents additional challenges for retrieval systems, requiring them to adapt to the dataset characteristic and specific balance of information.
Table 5: Positive/Non-negative rates (%) from human evaluation.
Naturalness Diversity Practicality
STaRK-Amazon 73.6 / 89.5 68.4 / 89.5 89.5 / 94.7
STaRK-MAG 94.7 / 100 73.7 / 84.2 68.4 / 84.2
STaRK-Prime 67.8 / 92.8 71.4 / 82.1 71.4 / 89.3
Average 78.7 / 94.1 71.0 / 85.3 76.4 / 89.4
Human evaluation. Moreover, we qualitatively assess sampled queries from our benchmark to determine if they are natural-sounding (resembling natural conversation), diverse (covering a wide range of question structures and complexity levels), and practical (relevant and useful in real-life situations), with participation from 63 individuals. The evaluation results are converted from a 5-point Likert-like scale to a positive/tie/negative scale for reporting. We report the positive and non-negative rates in Table 5. On average across the three datasets, 94.1%, 85.3%, and 89.4% of participants rated neutral or above in terms of naturalness, diversity, and practicality, respectively. The results justify the quality of our benchmark and highlight its potential for diverse and realistic retrieval tasks.
4 Experiments
In this section, we aim to explore the following questions:
  • Q1: How well do current retrieval approaches perform on our benchmark datasets?
  • Q2: What challenges are revealed by the experiments, and more importantly, what are the potential future directions for improvement?
4.1 Retrieval Models
Vector Similarity Search (VSS). VSS embeds both the query and the concatenated textual and relational information of each candidate entity. Then, the similarity score is computed based on cosine similarity between the query and candidate embeddings. We use the text-embedding-ada-002 model from OpenAI for generating the embeddings.
Multi-Vector Similarity Search (Multi-VSS). Multi-VSS represents candidate entities with multiple vectors, capturing detailed features for complex retrieval tasks. Here, we chunk and separately embed the textual and relational information of each candidate entity. The final score is aggregated from the similarities between the chunks and the query.
Dense Retriever. Dense Retriever finetunes a query encoder and a document encoder separately using the query answer pairs from the training dataset. We optimize the encodes using contrastive loss. Specifically, the positive pairs are constructed from the training questions and their ground truth answers, while we construct 20 hard negative answers for each query from the top false positive predictions from VSS. We use roberta-base as the base model and finetune the encodes over the entire training split for each dataset.
QAGNN [47]. QAGNN constructs a graph where nodes include entities found in the question or answer choices, incorporating their neighboring nodes as intermediate context. It extracts a subgraph from a larger knowledge graph, enabling a comprehensive understanding by leveraging the relational information from the knowledge graph. The approach integrates this with semantic embeddings from a language model, jointly modeling both relational and semantic information to enhance the question-answering modeling.
VSS + LLM Reranker [9, 53]. This method improves the precision of the top-$v$ results from VSS by reranking them using language models, taking advantage of their advanced language understanding capabilities. For this purpose, we employ two different language models: GPT-4-turbo (gpt-4-1106-preview) and Claude-v2. We set the value of $v=20$ for the reranking process to balance between precision and computational efficiency. Specifically, we construct a prompt that instructs the language models to assign a score between 0 and 1 to a node, based on its combined textual and relational information, with certain criteria provided for rating the node. We find that this approach, which involves directly scoring nodes, performs better than asking for strict satisfaction of conditions [53].
4.2 Evaluation Metrics
We use the following metrics for a holistically evaluating the model performance.
Hit@$k$. The Hit@$k$ metric assesses whether the correct item is among the top-$k$ results from the model. We used $k=1$ and $k=5$ for evaluation. At $k=1$, it evaluates the accuracy of the top recommendation; at $k=5$, it examines the model’s precision in a wider recommendation set. It’s a straightforward yet effective way to measure the immediate relevance of answers.
Recall@$k$. Recall@$k$ measures the proportion of relevant items in the top-$k$ results. In our study, $k=20$ is used, as the answer length of all of the queries in our benchmarks are equal or smaller then $20$. This metric offers insight into the model’s ability to identify all relevant items, particularly in scenarios where missing any could be critical.
Mean Reciprocal Rank (MRR). MRR is a statistic for evaluating the average effectiveness of a predictive model. It calculates the reciprocal of the rank at which the first relevant item appears in the list of predictions. MRR is particularly useful for understanding the model’s performance at presenting the correct item at an early stage, without the need for a predefined $k$. This metric emphasizes the importance of the rank of the first correct answer, which is crucial in many practical applications where the first correct answer is often the most impactful.
4.3 Analysis
Table 6: Main experimental results.
Dense Retriever QAGNN VSS Multi-VSS VSS+Claude2 VSS+GPT4
(roberta) (roberta) (ada-002) (ada-002) Reranker Reranker
STaRK-Amazon Hit@1 0.1529 0.2656 0.3916 0.4007 0.4266 0.4479
Hit@5 0.4793 0.5001 0.6273 0.6498 0.6746 0.7117
Recall@20 0.4449 0.5205 0.5329 0.5512 0.5376 0.5535
MRR 0.3020 0.3775 0.5035 0.5155 0.5329 0.5569
STaRK-MAG Hit@1 0.1051 0.1288 0.2908 0.2592 0.3202 0.4090
Hit@5 0.3523 0.3901 0.4961 0.5043 0.5334 0.5818
Recall@20 0.4211 0.4697 0.4836 0.5080 0.4834 0.4860
MRR 0.2134 0.2912 0.3862 0.3694 0.4129 0.4900
STaRK-Prime Hit@1 0.0446 0.0885 0.1263 0.1510 0.1611 0.1828
Hit@5 0.2185 0.2135 0.3149 0.3356 0.3582 0.3728
Recall@20 0.3013 0.2963 0.3600 0.3805 0.3598 0.3405
MRR 0.1238 0.1473 0.2141 0.2349 0.2466 0.2655
Main results. In Table 6, we report the experiment results. Firstly, we notice a significant performance gap between models trained or fine-tuned on our training datasets, such as Dense Retriever and QAGNN, and models like VSS and Multi-VSS that use embeddings generated from text-embedding-ada-002. For the dense retriever, training the encoders proves challenging when handling both textual and relational information as text, leading to training challenges. The difficulty in training a dense retriever is exacerbated by the stark differences in length between long documents and short relational pairs, making it tough for the model to learn a new knowledge base’s structure from scratch during fine-tuning. Without the advantage of pre-trained embeddings, it also struggles to align with the textual requirements. For QAGNN, it faces its own training challenges due to its high computational demands during training. Finally, the superiority of VSS and Multi-VSS demonstrates the superior pre-trained embeddings obtained from a large corpus.
Further, the Multi-VSS strategy, which breaks down the documents into chunks and embeds them into embeddings separately to capture more fine-grained context, generally outperforms VSS. However, it falls short in the Hit@1 metric on the STaRK-MAG dataset, likely due to this dataset’s property for understanding more extended contexts or potential interaction or reasoning between different chunks.
Significantly, the methods combining VSS with an LLM reranker consistently outperform other approaches across most metrics. Notably, reranking the top 20 predictions with an LLM leads to significant improvements in Hit@$k$ and MRR metrics compared to using VSS alone. This improvement underscores the limitations of traditional embedding methods, which lack the advanced reasoning capabilities provided by LLMs. Additionally, using GPT-4 as a reranker tends to yield better results than using Claude, further demonstrating the effectiveness of better reasoning and context understanding in improving retrieval performance.
However, even for VSS+GPT4 Reranker, we observe insufficient performances, e.g., Hit@1 is only around 18% on STaRK-Prime and 41% on STaRK-MAG , meaning that for a significant portion of queries, the top-ranked answer is not the correct one. Moreover, Recall@20 metric values are lower than 60% across all of the datasets, which indicates the ranking results may not be comprehensive, potentially missing relevant answers in specific applications. Also, we notice that the MRR scores generally follow a similar trend as the Hit@1 metric. However, the MRR values remain relatively low, particularly for STaRK-Prime, indicating large space for further optimization in the ranking process.
Table 7: Latency (s) of the retrieval systems on STaRK.
Dense Retriever QAGNN VSS Multi-VSS VSS+Claude VSS+GPT4
STaRK-Amazon 2.34 2.32 5.71 4.87 27.24 24.76
STaRK-MAG 0.94 1.35 2.25 3.14 22.60 23.43
STaRK-Prime 0.92 1.29 0.54 0.90 29.14 26.97
Average 1.40 1.65 2.83 2.97 26.33 25.05
Retrieval Latency. Another aspect vital for the practical application of retrieval systems is their latency. As shown in Table 7, we test the latency of the retrieval methods on a single NVIDIA A100-SXM4-80GB. Specifically, the Dense Retriever and QAGNN models exhibit lower average latency, making them more suitable for time-sensitive applications. In contrast, the VSS and Multi-VSS models show moderate latency, but when combined with LLM rerankers, such as Claude and GPT-4, the latency significantly increases. Therefore, in the context of semi-structured retrieval tasks, the balance between enhanced accuracy and increased latency is crucial. Effective retrieval methods should consider such trade-offs, particularly when addressing complex queries involving reasoning over both textual and relational information.
Figure 6: A case study on STaRK-MAG, where VSS mistakenly ranks non-ground truth papers C and D higher due to the repeated key words in the relational information "cites paper." After reranking with Claude, it correctly prioritizes the ground truth papers A and B. This correction is attributed to the more accurate reasoning and analysis of the combined textual and relational information.
Case study. Finally, we highlight the importance of reasoning ability in achieving good retrieval results on our benchmark datasets. In Figure 6, we conduct a case study to compare the performance of VSS and VSS+Claude Reranker. Specifically, we examine a user query that requests papers from an institution on a particular topic. In this scenario, VSS fails to adequately address the relational requirement due to directly embedding the entire documents without detailed analysis. As a result, papers containing frequently repeated keywords, such as “nonlinear modeling” and “piezoelectric elements”, are mistakenly assigned high scores. However, the result improves significantly after the reranking with an LLM, where the LLM is equipped to reason the relationship between the query and the information contained in each paper, making the scores more accurately reflect the papers’ relevance to the query. This case study underscores the need for reasoning ability in grasping the complexities of queries. As embedding methods may capture some aspects of the query, they often fail to capture the detailed interaction between textual and relational information. Thus, integrating LLMs with advanced reasoning into the retrieval process can be crucial for the retrieval accuracy.
5 Conclusion
We introduce STaRK, the first benchmark designed to thoroughly evaluate the capability of retrieval systems driven by LLMs in handling semi-structured knowledge bases (SKBs). This benchmark features a diverse set of queries that are semi-structured and natural-sounding, requiring context-specific reasoning across various domains, thereby setting a new standard for assessing retrieval systems in the context of SKBs. We utilize public datasets to construct three SKBs and develop an automated, general pipeline to simulate user queries that mimic real-world scenarios. Our extensive experiments on STaRK reveal significant challenges for current models in handling both textual and relational information effectively and flexibly. Overall, STaRK offers valuable opportunities for future research to advance the field of complex and multimodal retrieval systems, where reducing retrieval latency and incorporating strong reasoning ability into the retrieval process are identified as two prospective future directions.
6 Acknowledgement
We thank group members in Jure Leskovec and James Zou’s lab for providing valuable suggestions. We also acknowledge the support of Amazon, DARPA under Nos. N660011924033 (MCS); NSF under Nos. OAC-1835598 (CINES), CCF-1918940 (Expeditions), DMS-2327709 (IHBEM); Stanford Data Applications Initiative, Wu Tsai Neurosciences Institute, Chan Zuckerberg Initiative, Genentech, GSK, Hitachi, Juniper Networks, KDDI, and UCB.
Appendix A Benchmark details
relational structure textual information
STaRK-Amazon Amazon Product Reviews Amazon Product Reviews
Amazon Question and Answer Data
STaRK-MAG ogbn-mag ogbn-papers100M, Microsoft Academic Graph
STaRK-Prime PrimeKG disease: Orphanet; drug: DrugBank; pathway: Reactome;
gene: Ensembl, NCBI Entrez, Uniprot, UCSC, CPDB
Table 8: Sources of relational structure and textual information of the benchmarks
A.1 STaRK-MAG
Relational query templates:
metapath multi-hop query template
(author → paper) "Can you list the papers authored by <author>?"
(paper → paper) "Which papers have been cited by the paper <paper>?"
(field_of_study → paper) "Can you provide a list of papers in the field of <field_of_study>?"
(institution → author → paper) "What papers have been published by researchers from <institution>?"
(paper → author → paper) "What papers have been published by researchers that are coauthors of <paper>?"
(paper → author → paper ← field_of_study ← paper) "Can you find papers that share a coauthor with <paper> and are also in the same field of study?"
(institution → author → paper ← field_of_study) "Are there any papers associated with <institution> and are in the field of <field_of_study>?"
For example, the metapath (field_of_study $\rightarrow$ paper) requires an initial field_of_study entity to be filled in the corresponding query template. For multi-hop metapaths, the last metapath (institution $\rightarrow$ author $\rightarrow$ paper $\leftarrow$ field_of_study) requires an institution entity and a field_of_study entity to initialize the query.
A.2 STaRK-Prime
Number of entities in each type.
#disease:            17,080
#gene/protein:       27,671
#molecular_function: 11,169
#drug:                7,957
#pathway:             2,516
#anatomy:            14,035
#effect/phenotype:   15,311
#biological_process: 28,642
#cellular_component:  4,176
#exposure:              818
Relational tuples.
[(anatomy, expression absent, gene/protein),
(anatomy, expression present, gene/protein),
(anatomy, parent-child, anatomy),
(biological_process, interacts with, exposure),
(biological_process, interacts with, gene/protein),
(biological_process, parent-child, biological_process),
(cellular_component, interacts with, exposure),
(cellular_component, interacts with, gene/protein),
(cellular_component, parent-child, cellular_component),
(disease, associated with, gene/protein),
(disease, contraindication, drug),
(disease, indication, drug),
(disease, linked to, exposure),
(disease, off-label use, drug),
(disease, parent-child, disease),
(disease, phenotype absent, effect/phenotype),
(disease, phenotype present, effect/phenotype),
(drug, carrier, gene/protein),
(drug, contraindication, disease),
(drug, enzyme, gene/protein),
(drug, indication, disease),
(drug, off-label use, disease),
(drug, side effect, effect/phenotype),
(drug, synergistic interaction, drug),
(drug, target, gene/protein),
(drug, transporter, gene/protein),
(effect/phenotype, associated with, gene/protein),
(effect/phenotype, parent-child, effect/phenotype),
(effect/phenotype, phenotype absent, disease),
(effect/phenotype, phenotype present, disease),
(effect/phenotype, side effect, drug),
(exposure, interacts with, biological_process),
(exposure, interacts with, cellular_component),
(exposure, interacts with, gene/protein),
(exposure, interacts with, molecular_function),
(exposure, linked to, disease),
(exposure, parent-child, exposure),
(gene/protein, associated with, disease),
(gene/protein, associated with, effect/phenotype),
(gene/protein, carrier, drug),
(gene/protein, enzyme, drug),
(gene/protein, expression absent, anatomy),
(gene/protein, expression present, anatomy),
(gene/protein, interacts with, biological_process),
(gene/protein, interacts with, cellular_component),
(gene/protein, interacts with, exposure),
(gene/protein, interacts with, molecular_function),
(gene/protein, interacts with, pathway),
(gene/protein, ppi, gene/protein),
(gene/protein, target, drug),
(gene/protein, transporter, drug),
(molecular_function, interacts with, exposure),
(molecular_function, interacts with, gene/protein),
(molecular_function, parent-child, molecular_function),
(pathway, interacts with, gene/protein),
(pathway, parent-child, pathway)]
Relational query templates.
(effect/phenotype → [phenotype absent] → disease ← [!indication] ← drug):
    "Find diseases with zero indication drug and are associated with <effect/phenotype>",
(drug → [contraindication] → disease ← [associated with] ← gene/protein):
    "Identify diseases associated with <gene/protein> and are contraindicated with <drug>",
(anatomy → [expression present] → gene/protein ← [expression absent] ← anatomy):
    "What gene or protein is expressed in <anatomy1> while is absent in <anatomy2>?",
(anatomy → [expression absent] → gene/protein ← [expression absent] ← anatomy):
    "What gene/protein is absent in both <anatomy1> and <anatomy2>?",
(drug → [carrier] → gene/protein ← [carrier] ← drug):
    "Which target genes are shared carriers between <drug1> and <drug2>?",
(anatomy → [expression present] → gene/protein → [target] → drug):
    "What is the drug that targets the genes or proteins which are expressed in <anatomy>?",
(drug → [side effect] → effect/phenotype → [side effect] → drug):
    "What drug has common side effects as <drug>?",
(drug → [carrier] → gene/protein → [carrier] → drug):
    "What is the drug that has common gene/protein carrier with <drug>?",
(anatomy → [expression present] → gene/protein → enzyme → drug):
    "What is the drug that some genes or proteins act as an enzyme upon, where the genes or proteins are expressed in <anatomy>?",
(cellular_component → [interacts with] → gene/protein → [carrier] → drug):
    "What is the drug carried by genes or proteins that interact with <cellular_component>?",
(molecular_function → [interacts with] → gene/protein → [target] → drug):
    "What drug targets the genes or proteins that interact with <molecular_function>?",
(effect/phenotype → [side effect] → drug → [synergistic interaction] → drug):
    "What drug has a synergistic interaction with the drug that has <effect/phenotype> as a side effect?",
(disease → [indication] → drug → [contraindication] → disease):
    "What disease is a contraindication for the drugs indicated for <disease>?",
(disease → [parent-child] → disease → [phenotype present] → effect/phenotype):
    "What effect or phenotype is present in the sub type of <disease>?",
(gene/protein → [transporter] → drug → [side effect] → effect/phenotype):
    "What effect or phenotype is a [side effect] of the drug transported by <gene/protein>?",
(drug → [transporter] → gene/protein → [interacts with] → exposure):
    "What exposure may affect <drug>s efficacy by acting on its transporter genes?",
(pathway → [interacts with] → gene/protein → [ppi] → gene/protein):
    "What gene/protein interacts with the gene/protein that related to <pathway>?",
(drug → [synergistic interaction] → drug → [transporter] → gene/protein):
    "What gene or protein transports the drugs that have a synergistic interaction with <drug>?",
(biological_process → [interacts with] → gene/protein → [interacts with] → biological_process):
    "What biological process has the common interactino pattern with gene or proteins as <biological_process>?",
(effect/phenotype → [associated with] → gene/protein → [interacts with] → biological_process):
    "What biological process interacts with the gene/protein associated with <effect/phenotype>?",
(drug → [transporter] → gene/protein → [expression present] → anatomy):
    "What anatomy expressesed by the gene/protein that affect the transporter of <drug>?",
(drug → [target] → gene/protein → [interacts with] → cellular_component):
    "What cellular component interacts with genes or proteins targeted by <drug>?",
(biological_process → [interacts with] → gene/protein → [expression absent] → anatomy):
    "What anatomy does not express the genes or proteins that interacts with <biological_process>?",
(effect/phenotype → [associated with] → gene/protein → [expression absent] → anatomy):
    "What anatomy does not express the genes or proteins associated with <effect/phenotype>?",
(drug → [indication] → disease → [indication] → drug) & (drug → [synergistic interaction] → drug):
    "Find drugs that has a synergistic interaction with <drug> and both are indicated for the same disease.",
(pathway → [interacts with] → gene/protein → [interacts with] → pathway) & (pathway → [parent-child] → pathway):
    "Find pathway that is related with <pathway> and both can [interacts with] the same gene/protein.",
(gene/protein → [associated with] → disease → [associated with] → gene/protein) & (gene/protein → [ppi] → gene/protein):
    "Find gene/protein that can interect with <gene/protein> and both are associated with the same disease.",
(gene/protein → [associated with] → effect/phenotype → [associated with] → gene/protein) & (gene/protein → [ppi] → gene/protein):
    "Find gene/protein that can interect with <gene/protein> and both are associated with the same effect/phenotype."
where []delimited-[][$\cdot$][ ⋅ ] denotes the relation type.
Appendix B Prompts
B.1 Extracting textual requirements
{mdframed}
[backgroundcolor=cyan!5, linecolor=black, linewidth=2pt, innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin=3pt, innerbottommargin=10pt]
Prompt for STaRK-Amazon: Textual requirement extraction
You are an intelligent assistant that extracts diverse positive requirements and negative perspectives for an Amazon product. I will give you the following information:
- product: <product name>
- dimensions: <product dimensions>
- weight: <product weight>
- description: <product description>
- features:
#1: <feature #1>
- reviews:
#1:
summary: <review summary>
text: <full review text>
#2:
- Q&A:
#1:
question: <product-related question>
answer: <answer to product-related question>
#2:
Based on the given product information, you need to (1) identify the products generic category, (2) list all of the negative perspectives and their sources, and (2) extract up to five hard and five soft requirements relevant to customers interests along with their sources. (1) For example, the products generic category can be "a chess book" or "a phone case for iphone 6", do not use the product name directly. (2) Negative perspectives are those that the product doesnt fulfill, which come from the negative reviews or Q&A. (3) For the requirements, you should only focus on the products advantages and positive perspects. Hard requirements mean that product must fulfil, such as size and functionality. Soft requirements are not as strictly defined but still desirable, such as a product is easy-to-use. For (2) and (3), each source is a composite of the key and index (if applicable) separated by "-", such as "description", "Q&A-#1". You should provide the response in a specific format as follows where "item" refers to the products generic category, e.g., "a chess book".
Response format:
"item": <the products generic category> ,
"negative": [[<source of negative perspective>, <negative perspective description>]],
"hard": [[<source of hard requirement>, <hard requirement description>], ...],
"soft": [[<source of soft requirement>, <soft requirement description>], ...]
Here is an example of the response:
"item": "a camping chair",
"negative": [["reviews-#3", "the chair is not sturdy enough"], ["Q&A-#1", "wrong color"]],
"hard": [["description", "has a breathable mesh back"], ["description", "the arm is adjustable"], ["dimensions", "more than 35 inches long"], ["features-#7", "with a arm rest cup holder"], ["Q&A-#4", "need to come with a carrying bag"]],
"soft": [["description", "suitable for outdoors"], ["features-#9", "compact and save space"], ["reviews-#6", "light and portable"]]
This is the information of the product that you need to write response for:
Information:
<product_doc>
Response:
{mdframed}
[backgroundcolor=pink!25, linecolor=black, linewidth=2pt, innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin=3pt, innerbottommargin=10pt]
Prompt for STaRK-MAG: Textual requirement extraction
You are a helpful assistant that helps me extract one short requirement (no more than 10 words) about a paper from the paper information that researchers might be interested in. The requirement can be about the paper content, publication date, publication venue, etc. The requirement should be general and not too specific. I will give you the paper information, and you should return a short phrase about the paper, starting with the paper...’. This is the paper information:
<doc_info>
Please only return the short and general requirement without additional comments.
{mdframed}
[backgroundcolor=lime!10, linecolor=black, linewidth=2pt, innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin=3pt, innerbottommargin=10pt]
Prompt for STaRK-Prime: Textual requirement extraction
You are a helpful assistant that helps me extract <n_properties> from a given <target> information that a <role> may be interested in.
<role_instruction>
Each property should be no more than 10 words and start with "the <target>". You should also include the source of each property as indicated in the paragraph names of the information, e.g., "details.mayo_symptoms", "details.summary", etc. You should return a list of properties and their sources following the format:
[["<short_property1>", "<source1>"], ["<short_property2>", "<source2>"], ...]
This is the information:
<doc_info>
Please provide only the list with <n_properties> in your response. Response:
According to the role assigned to simulate the query content, the <role_instruction> as shown below is filled in accordingly.
role role instruction
Doctor Doctors typically ask questions aimed at diagnosing and treating. Their questions tend to be direct and practical, focusing on aspects involving side effects, symptoms, and complications etc.
Medical scientist Medical scientists often ask questions that reflect the complexity and depth of the scientific inquiry in the medical field. Their questions tend to be detailed and specific, focusing on aspects such as: etiology and pathophysiology, genetic factors, association with pathway, protein, or molecular function.
Patient Patients typically don’t know the professional medical terminology. Their questions tend to be straightforward, focusing on practical concerns on the symptons, effects, and inheritance etc., instead of the detailed mechanisms, which may also include more context.
B.2 Combining relational and textual requirements
{mdframed}
[backgroundcolor=cyan!5, linecolor=black, linewidth=2pt, innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin=3pt, innerbottommargin=10pt]
Prompt for STaRK-Amazon: Fuse relational and textual requirements
You are an intelligent assistant that generates queries about an Amazon item. I will provide you with the item name, requirements, and its negative customer reviews. Your task is to create a natural-sounding customer query that leads to the item as the answer, using the requirements that are non-conflicting with the negative reviews, and provide the indices of the requirements used. For example:
Information:
- item: a soccer rebounder
- requirements:
#1: needs a heavy-duty 1-inch to 3-inch steel tube frame
#2: should be adjustable for practicing different skills
#3: should be durable
#4: usually be viewed together with <SKLZ Star-Kick Hands Free Solo Soccer Trainer>
- negative reviews:
#1: it was broken after a few uses
Response:
"index": [1, 2, 4],
"query": "Please recommend a soccer rebounder with a steel frame, about 2 inches thick, that can adapt to different skill levels. We had a blast using the <SKLZ Star-Kick Hands Free Solo Soccer Trainer> with my family, and Im on the lookout for something similar."
As the negative review indicates that the soccer rebouncer lacks durability, your query should only incorporate requirements #1, #2, and #4 while excluding #3. A requirement should only be excluded if it conflicts with negative feedback or is unlikely to align with customers interests. For relational requirements about another <product>, do not directly use "usually bought/viewed together with <product>" in the query. You must deduce the items relationship with <product> into substitute or complement, and create various user scenarios, such as the item should be compatible or used with <product> (for complements) or match in style with <product> (for substitute), to make the queries sound natural. Except for <product>, you should change the description but convey similar meanings. The query structure is completely flexible. Here is the information to generate the requirement indices and a natural-sounding query:
Information:
<product_req_and_neg_comments>
Response:
{mdframed}
[backgroundcolor=lime!10, linecolor=black, linewidth=2pt, innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin=3pt, innerbottommargin=10pt]
Prompt for STaRK-MAG: Fuse relational and textual requirements
You are a helpful assistant that helps me generate a new query by incorporating an additional requirement into a given query, and form a coherent and natural-sounding question.
This is the existing query:
<query>
This is the additional requirement:
<additional_textual_requirement>
You should be creative in combining the existing query and requirement, and flexible in structuring the new query, adding context as needed. Please return the new query without additional comments:
The prompt of a second-time rewrite by GPT-4 Turbo:
You are a helpful assistant that helps make a researchers query about a paper more natural-sounding, akin to the language used in ArXiv web searches. You should change the description but convey similar meanings. The query structure is completely flexible. The original query:
"<query>"
Please only output the new query without additional comments:
{mdframed}
[backgroundcolor=pink!25, linecolor=black, linewidth=2pt, innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin=3pt, innerbottommargin=10pt]
Prompt for STaRK-Prime: Fuse relational and textual requirements
You are a helpful assistant that helps me generate a natural-sounding and coherent query as if you were a <role>. The query should be created based on a list of requirements for searching <plural_target> in a database. I will provide you with the requirements in the following format:
[<requirement1>, <requirement2>, ...]
You should create the query based solely on the given requirements. Moreover, you should craft the query from the perspective of a <role>.
<role_instruction>
For example, a query from a <role> could be
"<example_query>"
You can be flexible in structuring the query and adding additional context. Ensure that the query uses different descriptions than the original property descriptions while retaining similar meanings. The query should sound concise and natural. These are the requirements:
<requirements>
Please create the query based on the given requirements and provide only the query without additional comments. Your response:
The prompt of a second-time rewrite by GPT-4 Turbo:
You are a helpful assistant that helps me rewrite a query that searches for <plural_target> from the perspective of a <role>. You should maintain the requirements from the original query and the characteristics of the <role>, while being creative and flexible in structuring the query. Ensure the revised query is concise and natural-sounding. Original query: "<query>". Please output only the rewritten query:
B.3 Filtering additional answers
{mdframed}
[backgroundcolor=cyan!5, linecolor=black, linewidth=2pt, innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin=3pt, innerbottommargin=10pt]
Prompt for STaRK-Amazon: Filtering additional answers
Filter products by general category
You are an intelligent assistant that identifies whether an Amazon product belongs to a given category. I will give you the product information. You should only answer yes / no in the response. For examples, the product <SKLZ Star-Kick Hands Free Solo Soccer Trainer> belongs to the category "soccer trainer" and the product <Test your Opening, Middlegame and Endgame Play - VOLUME 2> belongs to the category "a chess opening book", while <Baby Girls One-piece Shiny Athletic Leotard Ballet Tutu with Bow> doesnt belong to category "an adult tutu".
Information:
- product title: <<product_title>>
- product description: <product_description>
Does the product belong to "<target_category>"? Response (yes/no):
Filter products by requirements
You are a helpful assistant that helps me check whether an Amazon product satisfies the given requirements. I will provide you with the product information, which may include the product description, features, reviews, and Q&A from customers. Your task is to assess whether the product meets each requirement based on the provided information. If there is no information that supports the requirement, your response for that requirement is "NA". If there is relevant information that supports the requirement, your response for that requirement is the information source that fulfills the requirement. Each information source is a composite of the key and index (if applicable), separated by "-", such as "description", "features-#3", "Q&A-#1", "reviews-#2". If there are multiple sources,
Response:
1: "NA" or [the information sources that satisfy the requirement #1],
2: "NA" or [the information sources that satisfy the requirement #2],
Here is the product information:
<product_doc>
The requirements are as follows:
<customer_requirements>
Response:
{mdframed}
[backgroundcolor=lime!10, linecolor=black, linewidth=2pt, innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin=3pt, innerbottommargin=10pt]
Prompt for STaRK-MAG: Filtering additional answers
You are a helpful assistant that helps me verify whether a given <target_node_type> is subject to a requirement. I will provide you with the <target_node_type> information and the requirement, and you should return only a True or False value, indicating whether the <target_node_type> meets the requirement.
This is the <target_node_type> information:
<doc_info>
This is the requirement:
<additional_textual_requirement>
Please return only the boolean value without additional comments:
{mdframed}
[backgroundcolor=pink!25, linecolor=black, linewidth=2pt, innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin=3pt, innerbottommargin=10pt]
Prompt for STaRK-Prime: Filtering additional answers
You are a helpful assistant tasked with verifying whether a given <target> satisfies each of the provided requirements. I will give you the requirements in the following format:
{1: <requirement1>, 2: <requirement2>, ...}
When evidence in the <target> information confirms a requirement is met, cite the source, for example, details.mayo_symptoms’, details.summary’. If no direct evidence exists, indicate this with NA’. The output in JSON format should be as follows:
{1: NA or <source1>, 2: NA or <source2>, ...}
This is the <target> information:
<doc_info>
These are the requirements:
<requirements>
Please provide only the JSON in your response. Response:"

### v2

@@
 Abstract
-Answering real-world user queries, such as product search, often requires accurate retrieval of information from semi-structured knowledge bases or databases that involve blend of unstructured (e.g., textual descriptions of products) and structured (e.g., entity relations of products) information.
+Answering real-world complex queries, such as complex product search, often requires accurate retrieval from semi-structured knowledge bases that involve blend of unstructured (e.g., textual descriptions of products) and structured (e.g., entity relations of products) information.
 However, previous works have mostly studied textual and relational retrieval tasks as separate topics.
 To address the gap, we develop STaRK, a large-scale Semi-structure retrieval benchmark on Textual and Relational Knowledge Bases.
-We design a novel pipeline to synthesize natural and realistic user queries that integrate diverse relational information and complex textual properties, as well as their ground-truth answers.
-Moreover, we rigorously conduct human evaluation to validate the quality of our benchmark, which covers a variety of practical applications, including product recommendations, academic paper searches, and precision medicine inquiries.
-Our benchmark serves as a comprehensive testbed for evaluating the performance of retrieval systems, with an emphasis on retrieval approaches driven by large language models (LLMs).
-Our experiments suggest that the STaRK datasets present significant challenges to the current retrieval and LLM systems, indicating the demand for building more capable retrieval systems that can handle both textual and relational aspects.
-Figure 1: STaRK features queries on semi-structured knowledge bases (SKBs) with textual and relational knowledge, with node entities as ground truth answers.
-It simulates user interactions with a SKB to evaluate LLM retrieval systems’ performance in providing accurate responses.
-Example query Title of ground truth node(s)
+Our benchmark covers three domains/datasets: product search, academic paper search, and queries in precision medicine.
+We design a novel pipeline to synthesize realistic user queries that integrate diverse relational information and complex textual properties, together with their ground-truth answers (items).
+We conduct rigorous human evaluation to validate the quality of our synthesized queries.
+We further enhance the benchmark with high-quality human-generated queries to provide an authentic reference.
+STaRK serves as a comprehensive testbed for evaluating the performance of retrieval systems driven by large language models (LLMs).
+Our experiments suggest that STaRK presents significant challenges to the current retrieval and LLM systems, indicating the demand for building more capable retrieval systems.
+The benchmark data and code are available on https://github.com/snap-stanford/stark.
+Figure 1: STaRK features queries on Semi-structured Knowledge Base (SKB) with textual and relational knowledge, with node entities as ground-truth answers.
+STaRK consists of synthesized queries simulating user interactions with a SKB and human-generated queries which provide an authentic reference.
+It evaluates LLM retrieval systems’ performance in providing accurate responses.
+Example query Title of ground-truth items(s)
 STaRK-Amazon Looking for durable Dart World brand dart flights that <Amazon Standard Flights>
-resist easy tearing. Any recommendations? <Dart World Broken Glass Flight> (12 more)
+resist easy tearing. Any recommendations? <Dart World Broken Glass Flight> (+12 more)
 What are recommended scuba diving weights for experienced divers <Sea Pearls Vinyl Coated Lace Thru Weight>
@@
 1 Introduction
-Natural-language queries are the primary forms in human society for acquiring information, involving diverse knowledge in the real world [15, 23, 19].
-For example, users on e-commerce web search can express complex information needs by combining free-form elements or constraints, such as “Can you help me find a push-along tricycle from Radio Flyer that’s both fun and safe for my kid?”.
-Here, the underlying knowledge can be represented in semi-structured formats [30, 41, 33], referred to as semi-structured knowledge bases (SKBs), which integrate unstructured data, such as textual descriptions and natural language expressions (e.g., descriptions about a tricycle), with structured data, like entity interactions on knowledge graph or database (e.g., a tricycle with “belongs to” relation with brand Radio Flyer).
-This allows semi-structured knowledge bases to represent comprehensive knowledge in specific applications, making them indispensable in domains such as e-commerce [13], social media [27], and biomedicine [6, 16].
-In this work, we focus on semi-structured knowledge bases (SKBs) with texts and relation information, which are two prevalent modalities on the existing knowledge bases.
-In this context, as shown in Figure 1, information retrieval serves as a predominant step by identifying relevant data from the semi-structured knowledge bases, such as specific tricycles that match the request, which are then processed to generate accurate answers to user queries [35].
-The retrieval accuracy on semi-structured knowledge bases is crucial for enhancing user experience, supporting informed decision-making, and preventing hallucination.
-Limitations of existing works.
-However, previous works and benchmarks have primarily focused on either purely textual queries on unstructured knowledge [18, 12, 22, 25, 44] or structured SQL or knowledge graph queries [2, 48, 46, 37], which are inadequate to study the complexities of retrieval tasks involving semi-structured knowledge bases.
+Natural-language queries are the primary form of how humans acquire information [15, 24, 19].
+For example, users on e-commerce sites wish to express complex information needs by combining free-form elements and constraints, such as “Can you help me find a push-along tricycle from Radio Flyer that’s both fun and safe for my kid?
+in product search.
+Medical scientists may ask questions like “What disease is associated with the PNPLA8 gene and presents with hypotonia as a symptom?.
+Answering such queries is crucial for enhancing user experience, supporting informed decision-making, and preventing hallucination.
+To answer such queries the underlying knowledge can be represented in semi-structured knowledge bases (SKBs) [31, 43, 34], which integrate unstructured data, such as textual descriptions and natural language expressions (e.g., description of the tricycle), with structured data, like entity interactions on knowledge graph (e.g., a tricycle “brand” is Radio Flyer).
+This allows the SKBs to represent comprehensive knowledge in specific applications, making them indispensable in domains such as e-commerce [13], social media [28], and precision medicine [6, 16, 21].
+Limitations of prior works.
+Prior works and benchmarks have primarily focused on either purely textual queries on unstructured knowledge [18, 12, 23, 26, 46] or structured SQL or knowledge graph queries [2, 50, 48, 38], which are inadequate to study the complexities of retrieval tasks involving SKBs.
 Another line of works [10, 22, 48, 44, 45, 39] have explored the middle ground via multi-hop reasoning within single or across multiple documents.
@@
 It is also important to note that existing works focus mainly on general knowledge, e.g., from Wikipedia.
-In fact, the knowledge may also come from private sources, necessitating information retrieval systems adaptable to private semi-structured knowledge bases.
+In fact, the knowledge may commonly come from private sources, requiring information retrieval systems to operate on private SKBs.
 Therefore, there is a gap in our understanding of how current LLM retrieval systems handle the complex interplay between textual and relational requirements in queries that optionally involve private knowledge.
-Challenges.
-To address this gap, our goal is to develop a benchmark for LLM retrieval on SKBs.
-However, the lack of real queries on SKBs has become an obstacle.
-Although retrieval datasets with synthetic queries is a valuable supplement, accurately simulating user queries on SKBs is particularly difficult.
-This difficulty arises from the interdependence of textual and relational information, which can lead to challenges in precisely filtering through millions of candidates to construct the ground truth answers.
+Present work.
+To address this gap, we present a large-scale Semi-structure retrieval benchmark on Textual and Relational Knowledge Bases (STaRK) (Figure 1).
+The key technical challenge that we solve is how to accurately simulate user queries on SKBs.
+This difficulty arises from the interdependence of textual and relational information, which can lead to challenges in precisely filtering through millions of candidates to construct the ground-truth answers.
 Additionally, ensuring that queries are relevant to real-world applications and resembles real-world scenarios adds further complexity to the benchmarking process.
-Our contribution: STaRK.
-As shown in Figure 1, we present a large-scale Semi-structure retrieval benchmark on Textual and Relational Knowledge Bases (STaRK).
-We propose a novel pipeline that simulates user queries and constructs precise ground truth answers using three SKBs built from extensive texts and millions of entity relations from public sources.
+We develop a novel pipeline that simulates user queries and constructs precise ground truth answers using three SKBs built from extensive texts and millions of entity relations from public sources.
 We validate the quality of queries in our benchmark through detailed analysis and human evaluation, focusing on their naturalness, diversity, and practicality.
-With STaRK, we delve deeper into retrieval tasks on SKBs, evaluate the capability of current retrieval systems, and providing insights for future advancement.
-We point out the key features of STaRK as follows:
-  • Natural-sounding queries on semi-structured knowledge: As illustrated in Table 1, the queries in our benchmark are crafted to incorporate rich relational information and complex textual properties.
+Furthermore, we incorporate 274 human-generated queries into our benchmark to compare with synthesized queries and enrich the testing scenarios.
+With STaRK, we delve deeper into retrieval tasks on SKBs, evaluate the capability of current retrieval systems, and provide insights for future advancement.
+Key features of STaRK are:
+  • Natural-sounding queries on SKBs (Table 1): Queries in our benchmark are crafted to incorporate rich relational information and complex textual properties.
 Additionally, these queries closely mirror the types of questions users would naturally ask in real-life scenarios, e.g., with flexible query formats and possibly with extra contexts.
@@
 Such reasoning capabilities are crucial for accurately interpreting and responding to the nuanced requirements of each query.
-  • Diverse domains: Our benchmark spans a wide range of content and domains, reflecting the diverse nature of real-world information needs.
-It covers areas on product recommendation, academic paper search, and precision medicine inquiry.
+  • Diverse domains: Our benchmark spans a wide range of domains, reflecting the diverse nature of real-world information needs.
+It covers three different knowledge bases1footnotetext: 1 We provide an interactive interface for exploring the SKBs at https://huggingface.co/spaces/snap-stanford/SKB-Explorer with applications on product recommendation, academic paper search, and precision medicine inquiry.
 Therefore, STaRK provides a comprehensive evaluation of retrieval systems across various contexts and knowledge domains.
-Moreover, we perform extensive experiments on each benchmark dataset for the prevailing retrieval systems.
+We also perform extensive experiments on each benchmark dataset for the prevailing retrieval systems.
 We highlight research challenges especially on the capability to handle textual and relational requirements and the latency of retrieval systems on large-scale SKBs that involve million-scale entities or relations.
-Finally, we provide insights into future directions to build more capable retrieval systems.
-The benchmark datsets and code are available on https://github.com/snap-stanford/stark.
-2 Related Work
-We identify three primary categories on retrieval-oriented question answering (QA) datasets based on the knowledge sources they utilize: unstructured, structured, and semi-structured knowledge bases.
-We discuss the differences and limitations in the previous works that we seek to address.
-Unstructured QA Datasets.
-The line of work explores retrieving answers from a single document [32] or multiple document sources [40, 45, 21, 42, 10].
-For example, SQuAD [32] focuses on extracting answers from a single document, testing model’s ability to understand and retrieve information within a specific context.
-For multi-document retrieval, HotpotQA [45] demands reasoning over multiple documents and includes sentence-level supporting facts, enhancing the complexity of the retrieval task.
-TriviaQA [21] combines questions from trivia games with evidence documents, evaluating both retrieval and reasoning over multiple sources.
-Moreover, some works [29, 24] leverage generated results by search engines as ground truth answers or information sources for retrieval.
-However, unstructured QA datasets or knowledge bases often lack the depth of relational reasoning commonly required in answering complex user queries.
-Structured QA Datasets.
-Within this category, the datasets challenge models to retrieve answer from structured knowledge bases, either knowledge graphs [3, 49, 39, 4, 5, 11, 14] or tabular data [51, 50].
-For example, ComplexWebQuestions [39] challenge model’s abilities to decompose complex queries and interpret queries using entities and relationships defined in knowledge graphs.
-GraphQA [14] textualizes graphs by describing node attributes and edges with natural language, offering language model a way to handle KBQA.
-For datasets with tabular data, WikiSQL [51] and Spider [50] both address the task of converting natural language questions to SQL queries, with WikiSQL concentrating on single-table databases and Spider extending to complex multi-tables.
-While these datasets primarily focus on relational information, the lack of textual information limits questions within predefined relationships and entities, which constrains the breadth of available information, resulting in less comprehensive and nuanced responses compared to the extensive insights drawn from abundant textual data.
-Semi-Structured QA Datasets.
-Another line of work focuses on semi-structured QA with tabular and textual data.
-WikiTableQuestions [31] emphasizes comprehension of table structure and textual content, reflecting the semi-structured nature of many real-world datasets.
-TabFact [7] and HybridQA [8] combine textual and tabular data, requiring models to validate statements or answer questions using both data types.
-TabMCQ [20] introduces multiple-choice questions based on tables, adding complexity in understanding and reasoning over semi-structured data.
-However, different from this work, these existing datasets mainly focus on tables as structured sources, which may not incorporate rich relational information naturally exists between entities.
-Additionally, previous attempts to combine text and tabular data through external web links or additional text columns often result in data structures that are hard to navigate.
-Therefore, there is an emerging need to benchmark retrieval tasks on textual and relational knowledge base, offering more insights beyond the scope of structured tabular sources.
+Finally, we provide insights into future directions to build more capable retrieval systems over private knowledge bases.
 2 Benchmarking Retrieval Tasks over Textual and Relational Knowledge
+Figure 2: Demonstration of Semi-structured Knowledge Bases, where each knowledge base combines both textual and relational information in a complex way, making the retrieval tasks challenging.
 Table 2: Data statistics of our constructed semi-structured knowledge bases
@@
 types types degree
-STaRK-Amazon 2 3 3.0 1,032,407 3,886,603 592,067,882
-STaRK-MAG 4 4 10.6 1,872,968 19,919,698 212,602,571
-STaRK-Prime 10 18 62.6 129,375 8,100,498 31,844,769
-Figure 2: Demonstration of Semi-structured Knowledge Bases, where each knowledge base combines both textual and relational information in a complex way, making the retrieval tasks challenging.
-We develop three large-scale retrieval datasets on three semi-structure knowledge bases (SKBs).
-In Section 3.1, we introduce the schema and scale of the SKBs, which integrate entities with rich relational and textual information.
-In Section 3.2, we introduce the retrieval datasets and the key characteristics of the queries.
-Then, we highlight our novel dataset construction pipeline in Section 3.3, which simulates user queries involving textual and relational knowledge and automatically generates the ground truth answers in the retrieval datasets.
-Finally, we conduct data distribution analysis and human evaluation on the queries in Section 3.4.
-3.1 Semi-structured Knowledge Bases
+STaRK-Amazon 4 4 18.2 1,035,542 9,443,802 592,067,882
+STaRK-MAG 4 4 43.5 1,872,968 39,802,116 212,602,571
+STaRK-Prime 10 18 125.2 129,375 8,100,498 31,844,769
+2.1 Problem Definition
+We are given a Semi-Structured Knowledge Base (SKB), which consists of a knowledge graph $G$ and a collection of free text documents $D$.
+Formally, let $G=(V,E)$ be the knowledge graph, where $V$ is the set of nodes and $E\subseteq V\times V$ is the set of edges representing relationships between nodes.
+$D=\bigcup_{i\in V}D_{i}$ be the collection of free-form text documents associated with the nodes, where $D_{i}$ is the set of documents associated with node $i$.
+For example, the product knowledge graph in e-commerce can capture relationships between products and brands/colors/categories, and the corresponding text documents include product descriptions, reviews, etc..
+We define the tasks on our benchmark datasets as follows: Given the knowledge graph $G=(V,E)$, a collection of free text documents $D$, and a query $Q$, the output is a set of nodes $A\subseteq V$ such that for each node $i\in A$, it satisfies the relational requirements imposed by the structure of $G$ as specified in $Q$, and the associated documents $D_{i}$ satisfy the textual requirements specified in $Q$.
+2.2 Semi-structured Knowledge Bases (SKBs)
 As shown in Figure 2, we construct three SKBs from the relational structure between entities and the textual information associated with a subset of the entities.
@@
 This knowledge base is derived from the Sports and Outdoors category from Amazon Product Reviews [13] and Amazon Question and Answer Data [29].
-The knowledge base features two entity types: product and brand, and three types of relational information: also_bought, also_viewed between product entities, and has_brand relation between product and brand entities.
-In total, it comprises around 1.03M entities (product entities: 0.96M, brand entities: 0.07M) and 3.9M relations (also_bought: 1.7M, also_viewed: 1.3M, has_a_brand: 0.9M).
+The knowledge base features four entity types: product, brand, color, and category, and four types of relational information: also_bought, also_viewed between product entities, has_brand, has_color, has_category relations between product and brand, color, and category entities, respectively.
+In total, it comprises around 1.0M entities (product entities: 0.9M, brand entities: 0.1M, category entities: 1.4k, color entities: 1.7k) and 9.4M relations (also_bought: 2.8M, also_viewed: 1.9M, has_brand: 1.7M, has_category: 2.3M, has_color: 0.6M).
 We obtain the textual information by combining the meta data from Amazon Product Reviews with the customer Q&A records from Amazon Question and Answer Data.
 This provides a rich amount of textual data, including product titles, descriptions, prices, customer reviews, and related customer Q&A for each product.
-For brand entities, we extract brand titles as the textual attribute.
-Especially, Amazon SKB features an extensive textual data, which is largely contributed from customer reviews and Q&A.
+For other entities, we extract their names or titles as the textual attribute.
+Amazon SKB features an extensive textual data, which is largely contributed from customer reviews and Q&A.
 MAG Semi-structured Knowledge Base.
 This knowledge base is constructed from obgn-MAG [17], obgn-papers100M [17], and Microsoft Academic Graph (version 2019-03-22) [43, 37].
-The knowledge base contains around 1.9M entities under four entity types (author: 1.1M, paper: 0.7M, institution: 9K, field_of_study: 0.06M) and 20M relations under four relation types (author_writes_paper: 7.9M, paper_has_field_of_study: 7.2M, paper_cites_paper: 4.9M, author_affiliated_with_institution: 1.0M).
+The knowledge base contains around 1.9M entities under four entity types (author: 1.1M, paper: 0.7M, institution: 8.7K, field_of_study: 59.5k) and 39.8M relations under four relation types (author_writes_paper: 13.5M, paper_has_field_of_study: 14.5M, paper_cites_paper: 9.7M, author_affiliated_with_institution: 2.0M).
 We construct the knowledge base by selecting papers1 shared between obgn-MAG and obgn-papers100M.
@@
 We leverage the existing knowledge graph PrimeKG [6] which contains ten entity types including disease, drug, gene/protein, and eighteen relation types, such as associated_with, synergistic_interaction, indication, expression_present.
-The entity count in our knowledge base is approximately 129K, with around 8M relations.
-The details on entity numbers and relation tuples are available in Appendix A.2.
+The entity count in our knowledge base is approximately 129.3K, with around 8.1M relations.
+The details on entity numbers and relation tuples are available in Appendix B.3.
 Compared to the Amazon and MAG SKBs, Prime SKB is denser and features a greater variety of relation types.
 While PrimeKG already provides text information on disease and drug entities, including drug mechanisms and disease descriptions, we additionally integrate the textual details from multiple databases for gene/protein and pathway entities such as genomic position, gene activity summary and pathway orthologous event.
-We include the comprehensive public sources we used to construct the SKBs in Table 8 of Appendix A.
+We include the public sources for SKBs construction in Table 10 (Appendix B).
 footnotetext: 1 We filter out non-English language papers as we only considers single-lingual queries.
 2.3 Retrieval Tasks on Semi-structured Knowledge Bases
-Table 3: Dataset statistics on STaRK.
-#queries #queries w/ avg. #answers train / val / test
-multiple answers
-STaRK-Amazon 9,100 7,082 17.99 0.65 / 0.17 / 0.18
+Table 3: Statistics on the STaRK benchmark datasets.
+#queries #queries w/ average train / val / test
+multiple answers #answers
+Synthesized (Sec 2.4, 2.5) STaRK-Amazon 9,100 7,082 17.99 0.65 / 0.17 / 0.18
 STaRK-MAG 13,323 6,872 2.78 0.60 / 0.20 / 0.20
 STaRK-Prime 11,204 4,188 2.56 0.55 / 0.20 / 0.25
-We develop three novel retrieval datasets.
-Specifically, the task is to retrieve node entities over the SKBs given any input queries.
-To highlight, the queries feature the interplay and fusion between relational and textual knowledge.
+Human-generated (Sec 2.6) STaRK-Amazon 81 64 19.50 For testing only
+STaRK-MAG 84 34 3.26
+STaRK-Prime 109 43 2.61
+We develop three novel retrieval-based question-answering datasets.
+To highlight, the queries feature the synthesization between relational and textual knowledge.
 Moreover, to elevate their applicability in practical scenarios, these queries are designed to mimic real-life query patterns in terms of the natural-sounding property and flexible formats.
-Each query may have single answer or multiple answers, formulated as a subset of entities from the knowledge base, which requires the model to accurately identify the ground truth entities out of potentially millions in response to each query.
-In Table 1, we demonstrate example queries from these benchmarks.
 We detail our retrieval benchmarks’ scale in Table 3 and introduce the specifics of each retrieval dataset as follows:
 STaRK-Amazon.
-This dataset comprises 9,100 customer-focused queries aimed at product searches, with a notable 68% of these queries yielding multiple answers.
+This dataset comprises synthesized 9,100 customer-focused queries along with 81 human-generated queries for additional evaluation.
+The task aims at product searches, with a notable 68% of the synthesized queries yielding multiple answers.
 The dataset prioritizes customer-oriented criteria, highlighting textual elements such as product quality, functionality, and style.
-Additionally, it accentuates relational aspects, including brand and product connections (e.g., complementary or substitute items).
+Additionally, it accentuates relational aspects, including brand, category, color and product connections (e.g., complementary or substitute items).
 The queries are framed in detailed, conversation-like formats, enriching the context and enhancing the dataset’s relevance to real-world scenarios.
 STaRK-MAG.
-This dataset comprises 13,323 paper queries, with roughly half yielding multiple answers.
-Beyond the single-hop relational requirements, STaRK-MAG also emphasizes the fusion between the textual requirements with multi-hop queries.
+This dataset comprises 13,323 synthesized and 84 human-generated queries targeting at academic paper search.
+Beyond the single-hop relational requirements on STaRK-Amazon, STaRK-MAG also emphasizes the fusion between the textual requirements with multi-hop queries.
 For example, “Are there any papers from King’s College London” highlights the metapath (institution $\rightarrow$ author $\rightarrow$ paper) on the relational structure.
 We designed three single-hop and four multi-hop relational query templates, cf.
-Appendix A.1.
+Appendix B.2.
 The textual aspects of the queries diversifies, focusing on different elements such as the paper’s topic, methodology, or the advancements it introduces, sourced mainly from the abstracts.
 STaRK-Prime.
-This dataset, comprising 11,204 queries across all ten entity types, incorporates single-hop and multi-hop relational information similar to STaRK-MAG.
-We developed 28 multi-hop query templates, detailed in Appendix A.2, to cover various relation types and ensure their practical relevance.
+This dataset, comprising 11,204 synthesized queries and 109 human-generated queries across all ten entity types, incorporates single-hop and multi-hop relational information similar to STaRK-MAG.
+For synthesized queries, we developed 28 multi-hop query templates, detailed in Appendix B.3, to cover various relation types and ensure their practical relevance.
 For instance, the query “What is the drug that targets the genes or proteins expressed in <anatomy>?” serves applications in precision medicine and pharmacogenomics, aiding researchers and healthcare professionals in identifying drugs that act on genes or proteins associated with specific anatomical areas and enabling more targeted treatments.
@@
 For entities such as effect/phenotype, the queries rely solely on relational data due to limited textual information in the knowledge base.
-We divide each dataset into training, validation, and testing subsets, with the ratios detailed in Table 3.
+We divide the synthesized queries on each dataset into training, validation, and testing subsets, with the ratios detailed in Table 3, while using the human-generated queries as additional sets for evaluation.
 For the STaRK-Amazon dataset, we randomly allocate a subset of queries with 20 or fewer answers to the validation and testing sets, while the remaining queries are assigned to the training set.
 For the STaRK-MAG and STaRK-Prime, we use random splits to divide the data.
-3.3 Benchmark Construction
+2.4 Benchmark Construction: Synthesized Queries
 Figure 3: The process of constructing semi-structured retrieval datasets involves four main steps: 1) Sample Relational Requirement: Based on relational templates, sample a relational requirement on a SKB.
@@
 4) Construct Ground Truth Nodes: Check if nodes satisfy the textual properties using multiple language models to establish ground truth nodes.
-Here we detail the novel pipeline to synthesize the retrieval datasets present in Section 3.2.
-The core idea is to entangle relational information and textual properties, while constructing ground truth answers in an accurate and efficient way.
-As shown in Figure 3, the construction of the retrieval datasets generally involves four steps, and the specific processes vary depending on the characteristics of each dataset.
+We present the novel pipeline that synthesizes the large-scale retrieval datasets.
+The core idea is to entangle relational information and textual properties and accurately construct ground truth answers.
+The construction of the retrieval datasets (Figure 3) generally involves four steps, and the specific processes vary depending on the characteristics of each dataset.
 These steps are as follows:
-  • Sample Relational Requirements: Initially, we sample a relation template, such as “(a product) belongs to <brand>” and instantiate it into a relational requirement, e.g., “belongs to Radio Flyer” shown in the first step of Figure 3.
+  • 1) Sample Relational Requirements: Initially, we sample a relation template, such as “(a product) belongs to <brand>” and ground it into a relational requirement, e.g., “belongs to Radio Flyer” in Figure 3.
 Each relational requirement yields a set of candidate entities that meet the requirement, i.e., products belonging to Radio Flyer.
-Therefore, we obtain the relational requirements and the corresponding candidates.
-The relation templates for the three SKBs are available in Appendix A, including 28 relation templates for the Prime SKB.
+The relation templates for the three SKBs are constructed with expert/domain knowledge, which are available in Appendix B.
 Each relation template in the SKBs serves a distinct and practical purpose.
@@
 It aids in pinpointing medications aimed at specific genetic or protein targets within defined anatomical areas, thereby contributing to the development of targeted treatments.
-  • Extracting Textual Properties: Subsequently, we select one of the candidate nodes that satisfies the relational requirement, which is referred to as the gold answer.
-We extract textual properties that align with the interests of specific roles from the textual information of the gold answer.
-The roles could be customers, researchers, or medical scientists, differing across the SKBs.
-For example, in the second step of Figure 3, we extract the phrases of functionality and user experience from a Radio Flyer product, which come from different sources in the product document.
-We utilize GPT-3.5-turbo-16k for the STaRK-Amazon and Claude-v2 for the other two datasets to conduct this step, where the prompts used are included in Appendix B.
-  • Combining Textual and Relational Information: With the textual and relational requirements, we simulate queries by fusing them using LLMs.
-We conduct two-stage fusion using two different LLMs, i.e., Claude-v2 and GPT-4-Turbo, to avoid bias that might arise from relying on a single LLM and ensure a more diverse set of simulated queries.
-Specifically, this first-stage integration is guided by various criteria, such as ensuring a natural-sounding query, adhering to the style of ArXiv searches, or aligning with specific roles.
-Furthermore, the second stage instructs the LLM to enrich the context and rephrase the language, thereby posing a more demanding reasoning challenge in comprehending the requirements of the query.
-In our example, by combining relational information “belongs to Radio Flyer” and textual information “push-along tricycle” and “fun and safe for kids”, we arrive at the final query in the third step.
-  • Filtering Additional Answers: In the final step, we assess whether each remaining candidate, excluding the gold answer used for extracting textual properties, meets the additional textual requirement.
-For the verification, we employ three Claude models.
-Only candidates that pass the verification across all models are included in the final ground truth answer set.
-Therefore, this stringent verification ensures the accuracy of the nodes in the ground truth answers.
-During this process, we also evaluate the accuracy of the gold nodes that pass the verification criteria.
-The accuracy rates STaRK-Amazon, STaRK-MAG, and STaRK-Primeare 86.6%, 98.9%, and 92.3%, respectively, demonstrating the effectiveness of our filtering approach in maintaining high-quality ground truth answers.
-This dataset construction pipeline is automatic, efficient, and broadly applicable to the SKBs within our context. We include the complete prompts of the above steps in Appendix B.
-3.4 Benchmark Data Distribution Analysis and Human Evaluation
+  • 2) Extracting Textual Properties: The relational requirement yields a set of satisfying candidates.
+Subsequently, we select one of the candidate nodes, referred to as the gold answer, and use LLMs to extract textual properties of interest from the its document.
+We extract the textual properties that align with the interests of specific roles, such as customers, researchers, or medical scientists, according to the specific SKB.
+For example, in Figure 3, we extract the phrases of functionality and user experience from a Radio Flyer product from different locations of the product document.
+  • 3) Combining Textual and Relational Information: After obtaining the textual and relational requirements, we synthesize them into a query using LLMs.
+We conduct two-stage synthesization using two different LLMs to avoid bias that might arise from relying on a single LLM and ensure a more diverse set of simulated queries.
+Specifically, this first-stage synthesization is guided by various criteria, such as natural sounding, adhering to the style of ArXiv searches, or aligning with specific roles.
+Furthermore, the second stage instructs the LLM to enrich the context and rephrase the language, demanding reasoning abilities in comprehending the requirements of the query.
+  • 4) Filtering Additional Answers: Finally, we employ multiple LLMs to verify if each remaining candidate, excluding the gold answer used for extracting textual properties, meets the additional textual requirement.
+Only candidates that pass the verification across all LLMs are included in the final ground truth answer set.
+This stringent verification ensures the accuracy of the nodes in the ground truth answers.
+To validate the precision of this filtering process, we compute the average precision of each gold answer passes the verification.
+The precision rates on STaRK-Amazon, STaRK-MAG, and STaRK-Prime are 86.6%, 98.9%, and 92.3%, respectively, demonstrating the effectiveness of our filtering approach in maintaining high-quality ground truth answers.
+This dataset construction pipeline is automatic, efficient, and broadly applicable to the SKBs in our formulation. We include all of the prompts and the LLMs versions in the above steps in Appendix C.
+2.5 Synthesized Data Distribution Analysis and Human Evaluation
 Figure 4: Distribution of query and answer lengths on STaRK datasets.
-Table 4: Query diversity measurement on STaRK.
+Table 4: Query diversity measurement on STaRK. See Appendix A for the metric definition.
 Shannon Type-Token
@@
 While TTR is more sensitive to the document length, the TTR values consistently demonstrate a steady presence of unique words relative to the large length in our query document.
-  • Ratio of Relational vs.
-Textual Information.
+  • Ratio of Relational vs.Textual Information.
 A key feature of our benchmark dataset is the composition of textual and relational information.
-Therefore, it is crucial to understand the proportionality between these two types of information.
-We calculate the ratio of the length of relational requirements to the length of textual requirements for each query and then average these ratios across each dataset.
+Therefore, it is crucial to understand the proportionality between these two types of information in the final queries.
+We calculate the ratio of relational to textual requirements for each query by comparing the word count of relational information to the word count of extracted textual properties, both of which are meta data in the constructing process.
+We then average these ratios across each dataset.
     Intuitively, the ratio serves as an approximation of the relative distribution based on the length of requirements, and does not directly reflect the importance of each type of information in determining the final answers.
@@
 The results justify the quality of our benchmark and highlight its potential for diverse and realistic retrieval tasks.
+2.6 Benchmark Construction: Human-Generated Queries
+To enhance the practical relevance of our benchmark dataset, we engaged 31 participants, 22 of whom are native English speakers, to generate a total of 274 queries across three SKBs.
+The participants were given a list of entity IDs, randomly selected from the entire entity set.
+They followed detailed instructions (see Appendix D) using our interactive platform to explore the SKB data.
+Subsequently, we manually filtered the ground truth answers, ensuring their accuracy through human verification.
+We then analyze the commonality and difference between synthesized and human-generated queries.
+Table 6: Comparison of Human-generated and Synthesized Queries
+Query Type Human-generated Query Synthesized Query
+Short and Direct Red sweatshirt for proud Montreal Canadiens Suggestions for a Suunto bike mount?
+Specific Author & Field Find me papers that discuss improving condenser performance authored by Stojan Hrnjak Show me papers by Seung-Hyeok Kye that discuss separability criteria.
+Complex Context Help me.
+I am trying to diagnose a patient with persistent joint pain, and I suspect a condition where the bone is dying due to compromised blood supply, often linked to factors like steroid use, alcohol abuse, or underlying diseases - what’s the name of this sneaky bone-killing culprit?
+I’m experiencing joint pain accompanied by swelling, fever… I’m concerned about medications aggravating my fuzzy eyesight and potential blood clotting complications.
+Could you recommend treatments effective for my symptoms while minimizing these specific side effects?
+Commonality.
+We present the statistics of the human-generated datasets in Table 3.
+The synthesized and human-generated queries exhibit a similar level of ambiguity, resulting in a comparable number of answers.
+Moreover, we find that most styles and formats of human-generated queries are represented in our synthesized query dataset.
+In Table 6, we demonstrate three pairs of human-generated and synthesized queries that display similarities such as short and straightforward product queries, inquiries targeting specific authors and fields, and complex contextual queries in STaRK-Prime.
+Difference.
+We find that human-generated queries often exhibit more unique expressions compared to synthesized queries, such as "Give me a fat cross and road tire that works with my Diamondback bicycle tube" and "this sneaky bone-killing culprit" (see the last row of Table 6).
+This discovery suggests a future direction for our benchmark to incorporate modern and dynamic language nuances.
 3 Experiments
@@
 Dense Retriever finetunes a query encoder and a document encoder separately using the query answer pairs from the training dataset.
-We optimize the encodes using contrastive loss.
+We optimize the encoder using contrastive loss.
 Specifically, the positive pairs are constructed from the training questions and their ground truth answers, while we construct 20 hard negative answers for each query from the top false positive predictions from VSS.
-We use roberta-base as the base model and finetune the encodes over the entire training split for each dataset.
+We use roberta-base as the base model and finetune the encoder over the entire training split for each dataset.
 QAGNN [49].
@@
 This method improves the precision of the top-$v$ results from VSS by reranking them using language models, taking advantage of their advanced language understanding capabilities.
-For this purpose, we employ two different language models: GPT-4-turbo (gpt-4-1106-preview) and Claude-v2.
-We set the value of $v=20$ for the reranking process to balance between precision and computational efficiency.
+For this purpose, we employ two different language models: GPT-4-turbo (gpt-4-1106-preview) and Claude3 (claude-3-opus).
+We set $v=20$ for synthesize queries and $v=10$ for human-generated queries to balance between precision and computational efficiency.
 Specifically, we construct a prompt that instructs the language models to assign a score between 0 and 1 to a node, based on its combined textual and relational information, with certain criteria provided for rating the node.
-We find that this approach, which involves directly scoring nodes, performs better than asking for strict satisfaction of conditions [53].
+Due to the high retrieval cost, we randomly sample 10% of the testing queries on each dataset for evaluation.
 3.2 Evaluation Metrics
@@
 Recall@$k$ measures the proportion of relevant items in the top-$k$ results.
-In our study, $k=20$ is used, as the answer length of all of the queries in our benchmarks are equal or smaller then $20$.
+For synthesized queries, $k=20$ is used, as the answer length of all of the queries in our benchmarks are equal or smaller then $20$.
+For human-generated queries, Recall@20 is not reported as the number of responses in the human-generated subset may exceed 20.
 This metric offers insight into the model’s ability to identify all relevant items, particularly in scenarios where missing any could be critical.
@@
 3.3 Analysis
-Table 6: Main experimental results.
-Dense Retriever QAGNN VSS Multi-VSS VSS+Claude2 VSS+GPT4
+Table 7: Testing results on STaRK-Syn(thesized). (*) indicates partial evaluation (10% of the testing queries) due to the high latency and cost of the methods.
+Dense Retriever QAGNN VSS Multi-VSS VSS+Claude3* VSS+GPT4*
 (roberta) (roberta) (ada-002) (ada-002) Reranker Reranker
-STaRK-Amazon Hit@1 0.1529 0.2656 0.3916 0.4007 0.4266 0.4479
-Hit@5 0.4793 0.5001 0.6273 0.6498 0.6746 0.7117
-Recall@20 0.4449 0.5205 0.5329 0.5512 0.5376 0.5535
-MRR 0.3020 0.3775 0.5035 0.5155 0.5329 0.5569
-STaRK-MAG Hit@1 0.1051 0.1288 0.2908 0.2592 0.3202 0.4090
-Hit@5 0.3523 0.3901 0.4961 0.5043 0.5334 0.5818
-Recall@20 0.4211 0.4697 0.4836 0.5080 0.4834 0.4860
-MRR 0.2134 0.2912 0.3862 0.3694 0.4129 0.4900
-STaRK-Prime Hit@1 0.0446 0.0885 0.1263 0.1510 0.1611 0.1828
-Hit@5 0.2185 0.2135 0.3149 0.3356 0.3582 0.3728
-Recall@20 0.3013 0.2963 0.3600 0.3805 0.3598 0.3405
-MRR 0.1238 0.1473 0.2141 0.2349 0.2466 0.2655
-Main results.
-In Table 6, we report the experiment results.
+STaRK-Amazon Hit@1 0.1529 0.2656 0.3916 0.4007 0.4549 0.4479
+Hit@5 0.4793 0.5001 0.6273 0.6498 0.7113 0.7117
+Recall@20 0.4449 0.5205 0.5329 0.5512 0.5377 0.5535
+MRR 0.3020 0.3775 0.5035 0.5155 0.5591 0.5569
+STaRK-MAG Hit@1 0.1051 0.1288 0.2908 0.2592 0.3654 0.4090
+Hit@5 0.3523 0.3901 0.4961 0.5043 0.5317 0.5818
+Recall@20 0.4211 0.4697 0.4836 0.5080 0.4836 0.4860
+MRR 0.2134 0.2912 0.3862 0.3694 0.4415 0.4900
+STaRK-Prime Hit@1 0.0446 0.0885 0.1263 0.1510 0.1779 0.1828
+Hit@5 0.2185 0.2135 0.3149 0.3356 0.3690 0.3728
+Recall@20 0.3013 0.2963 0.3600 0.3805 0.3557 0.3405
+MRR 0.1238 0.1473 0.2141 0.2349 0.2627 0.2655
+Table 8: Testing results on STaRK-Human(-Generated).
+VSS Multi-VSS VSS+Claude3 VSS+GPT4
+(ada-002) (ada-002) Reranker Reranker
+STaRK-Amazon Hit@1 0.3950 0.4691 0.5602 0.5432
+Hit@5 0.6419 0.7284 0.7284 0.7284
+MRR 0.5265 0.5874 0.6439 0.6274
+STaRK-MAG Hit@1 0.2857 0.2381 0.3810 0.3690
+Hit@5 0.4167 0.4167 0.4643 0.4523
+MRR 0.3581 0.3143 0.4243 0.4031
+STaRK-Prime Hit@1 0.2120 0.2567 0.2844 0.2844
+Hit@5 0.4037 0.4037 0.4771 0.4863
+MRR 0.2984 0.3377 0.3623 0.3617
+Results on Synthesized Queries.
+Table 7 shows the experiment results on the synthesized testing datasets.
 Firstly, we notice a significant performance gap between models trained or fine-tuned on our training datasets, such as Dense Retriever and QAGNN, and models like VSS and Multi-VSS that use embeddings generated from text-embedding-ada-002.
 For the dense retriever, training the encoders proves challenging when handling both textual and relational information as text, leading to training challenges.
-The difficulty in training a dense retriever is exacerbated by the stark differences in length between long documents and short relational pairs, making it tough for the model to learn a new knowledge base’s structure from scratch during fine-tuning.
-Without the advantage of pre-trained embeddings, it also struggles to align with the textual requirements.
-For QAGNN, it faces its own training challenges due to its high computational demands during training.
-Finally, the superiority of VSS and Multi-VSS demonstrates the superior pre-trained embeddings obtained from a large corpus.
+For QAGNN, it training challenges due to its high computational demands during training.
+Finally, the performance of VSS and Multi-VSS demonstrates the superior pre-trained embeddings obtained from a large corpus.
 Further, the Multi-VSS strategy, which breaks down the documents into chunks and embeds them into embeddings separately to capture more fine-grained context, generally outperforms VSS.
@@
 This improvement underscores the limitations of traditional embedding methods, which lack the advanced reasoning capabilities provided by LLMs.
-Additionally, using GPT-4 as a reranker tends to yield better results than using Claude, further demonstrating the effectiveness of better reasoning and context understanding in improving retrieval performance.
 However, even for VSS+GPT4 Reranker, we observe insufficient performances, e.g., Hit@1 is only around 18% on STaRK-Prime and 41% on STaRK-MAG , meaning that for a significant portion of queries, the top-ranked answer is not the correct one.
@@
 However, the MRR values remain relatively low, particularly for STaRK-Prime, indicating large space for further optimization in the ranking process.
+Results on human-generated dataset.
+We further evaluate the well-performing methods on human-generated queries.
+Table 8 presents the testing results on the human-generated dataset.
+Reranker methods continue to outperform other methods, confirming their superior reasoning and context understanding capabilities.
+Compared to synthesized datasets, the performance on human-generated queries shows highly correlated trends but generally higher metrics.
+This suggests that the synthesized datasets could be more difficult than the human-generated ones, offering challenging tasks and shedding light on interesting aspects of complex queries.
 Table 9: Latency (s) of the retrieval systems on STaRK.
@@
 Finally, we highlight the importance of reasoning ability in achieving good retrieval results on our benchmark datasets.
-In Figure 6, we conduct a case study to compare the performance of VSS and VSS+Claude Reranker.
+In Figure 6, we conduct a case study to compare the performance of VSS and VSS+Claude3 Reranker.
 Specifically, we examine a user query that requests papers from an institution on a particular topic.
@@
 Thus, integrating LLMs with advanced reasoning into the retrieval process can be crucial for the retrieval accuracy.
+4 Related Work
+We identify three primary categories on retrieval-oriented question answering (QA) datasets based on the knowledge sources they utilize: unstructured, structured, and semi-structured knowledge bases.
+We discuss the differences and limitations in the previous works that we seek to address.
+Unstructured QA Datasets.
+The line of work explores retrieving answers from a single document [33] or multiple document sources [42, 47, 22, 44, 10].
+For example, SQuAD [33] focuses on extracting answers from a single document, testing model’s ability to understand and retrieve information within a specific context.
+For multi-document retrieval, HotpotQA [47] demands reasoning over multiple documents and includes sentence-level supporting facts, enhancing the complexity of the retrieval task.
+TriviaQA [22] combines questions from trivia games with evidence documents, evaluating both retrieval and reasoning over multiple sources.
+Moreover, some works [30, 25] leverage generated results by search engines as ground truth answers or information sources for retrieval.
+However, unstructured QA datasets or knowledge bases often lack the depth of relational reasoning commonly required in answering complex user queries.
+In contrast, STaRK contains queries demanding multi-hop relational reasoning to challenge model’s ability of handling structured information.
+Structured QA Datasets.
+Within this category, the datasets challenge models to retrieve answers from structured knowledge bases, either knowledge graphs [3, 51, 40, 4, 5, 11, 14] or tabular data [53, 52].
+For example, ComplexWebQuestions [40] challenge models’ abilities to decompose complex queries and interpret queries using entities and relationships defined in knowledge graphs.
+GraphQA [14] textualizes graphs by describing node attributes and edges with natural language, offering language models a way to handle KBQA.
+For datasets with tabular data, WikiSQL [53] and Spider [52] both address the task of converting natural language questions to SQL queries, with WikiSQL concentrating on single-table databases and Spider extending to complex multi-tables.
+While these datasets primarily focus on relational information, the lack of textual information limits questions within predefined relationships and entities, which constrains the breadth of available information.
+STaRK extends this line of work by integrating textual information into knowledge bases and queries, resulting in more comprehensive and nuanced responses with extensive insights drawn from abundant textual data.
+Semi-Structured QA Datasets.
+Another line of work focuses on semi-structured QA with tabular and textual data.
+WikiTableQuestions [32] emphasizes comprehension of table structure and textual content, reflecting the semi-structured nature of many real-world datasets.
+TabFact [7] and HybridQA [8] combine textual and tabular data, requiring models to validate statements or answer questions using both data types.
+TabMCQ [20] introduces multiple-choice questions based on tables, adding complexity in understanding and reasoning over semi-structured data.
+However, different from this work, these existing datasets mainly focus on tables as structured sources, which may not incorporate rich relational information that naturally exists between entities.
+Additionally, previous attempts to combine text and tabular data through external web links or additional text columns often result in data structures that are hard to navigate.
+Therefore, STARK goes beyond these limitations by integrating both relational and textual data in a more cohesive manner, allowing for seamless navigation and retrieval across complex semi-structured knowledge bases.
+This integration supports richer relational reasoning and more comprehensive textual data handling, making STARK a more effective benchmark for evaluating the capabilities of retrieval systems in scenarios that closely mirror real-world information needs.
 5 Conclusion
@@
 We utilize public datasets to construct three SKBs and develop an automated, general pipeline to simulate user queries that mimic real-world scenarios.
+Moreover, we augment our datasets with human-generated queries and provide reference to the synthesized queries.
 Our extensive experiments on STaRK reveal significant challenges for current models in handling both textual and relational information effectively and flexibly.
@@
 6 Acknowledgement
-We thank group members in Jure Leskovec and James Zou’s lab for providing valuable suggestions.
-We also acknowledge the support of Amazon, DARPA under Nos.
+We thank group members group members in Leskovec and Zou labs for providing valuable suggestions and conducting benchmark construction.
+We express our gratitude to the following individuals for their assistance in generating the human-generated queries (ordered by last name):
+Michael Bereket, Charlotte Bunne, Yiqun Chen, Ian Covert, Alejandro Dobles, Teddy Ganea, Bryan He, Mika Sarkin Jain, Weixin Liang, Gavin Li, Jiayi Li, Sheng Liu, Michael Moor, Hamed Nilforoshan, Rishi Puri, Rishabh Ranjan, Yanay Rosen, Yangyi Shen, Jake Silberg, Elana Simon, Rok Sosic, Kyle Swanson, Nitya Thakkar, Rahul Thapa, Kevin Wu, Eric Wu, Kailas Vodrahalli.
+We especially thank Gavin Li at Stanford University and Zhanghan Wang at New York University for helping build the interactive interface for our SKBs.
+We also gratefully acknowledge the support of DARPA under Nos.
 N660011924033 (MCS); NSF under Nos.
-OAC-1835598 (CINES), CCF-1918940 (Expeditions), DMS-2327709 (IHBEM); Stanford Data Applications Initiative, Wu Tsai Neurosciences Institute, Chan Zuckerberg Initiative, Genentech, GSK, Hitachi, Juniper Networks, KDDI, and UCB.
-Appendix A Benchmark details
+OAC-1835598 (CINES), CCF-1918940 (Expeditions), DMS-2327709 (IHBEM); Stanford Data Applications Initiative, Wu Tsai Neurosciences Institute, Stanford Institute for Human-Centered AI, Chan Zuckerberg Initiative, Amazon, Genentech, GSK, Hitachi, SAP, and UCB.
+Appendix A Mathematical Definitions of Shannon Entropy and Type-Token Ratio
+Shannon Entropy. Shannon Entropy is a measure of the uncertainty in a set of possible outcomes, quantifying the amount of information or disorder within a dataset. It is defined as follows:
+$H(X)=-\sum_{i=1}^{n}p(x_{i})\log p(x_{i})$
+where $X$ is the set of possible outcomes, $p(x_{i})$ is the probability of occurrence of the outcome $x_{i}$, and $n$ is the total number of unique outcomes.
+Higher entropy values indicate greater diversity in the distribution of outcomes.[35]
+Type-Token Ratio (TTR).
+The Type-Token Ratio is a measure of lexical diversity, calculated as the ratio of the number of unique words (types) to the total number of words (tokens) in a text.
+It is defined as follows:
+$\text{TTR}=\frac{V}{N}$
+where $V$ is the number of unique words and $N$ is the total number of words in the text. Higher TTR values indicate a higher proportion of unique words, reflecting greater lexical diversity. [41]
+Appendix B Benchmark details
 relational structure textual information
@@
 Table 10: Sources of relational structure and textual information of the benchmarks
-A.1 STaRK-MAG
+B.1 STaRK-Amazon
+Basic query templates:
+metapath Query template
+(brand → product) "Can you list the products made by <brand>?"
+(product → product) "Which products are similar to <product>?"
+(color → product) "Can you provide a list of products that are available in <color>?"
+(category → product) "What products are available in the <category> category?"
+The final relational template can be composed of multiple basic templates. For example, ‘(color → product ← brand)’ represents one such combination.
+B.2 STaRK-MAG
 Relational query templates:
@@
 For multi-hop metapaths, the last metapath (institution $\rightarrow$ author $\rightarrow$ paper $\leftarrow$ field_of_study) requires an institution entity and a field_of_study entity to initialize the query.
-A.2 STaRK-Prime
+B.3 STaRK-Prime
 Number of entities in each type.
@@
 #exposure:              818
-Relational tuples.
-[(anatomy, expression absent, gene/protein),
-(anatomy, expression present, gene/protein),
-(anatomy, parent-child, anatomy),
-(biological_process, interacts with, exposure),
-(biological_process, interacts with, gene/protein),
-(biological_process, parent-child, biological_process),
-(cellular_component, interacts with, exposure),
-(cellular_component, interacts with, gene/protein),
-(cellular_component, parent-child, cellular_component),
-(disease, associated with, gene/protein),
-(disease, contraindication, drug),
-(disease, indication, drug),
-(disease, linked to, exposure),
-(disease, off-label use, drug),
-(disease, parent-child, disease),
-(disease, phenotype absent, effect/phenotype),
-(disease, phenotype present, effect/phenotype),
-(drug, carrier, gene/protein),
-(drug, contraindication, disease),
-(drug, enzyme, gene/protein),
-(drug, indication, disease),
-(drug, off-label use, disease),
-(drug, side effect, effect/phenotype),
-(drug, synergistic interaction, drug),
-(drug, target, gene/protein),
-(drug, transporter, gene/protein),
-(effect/phenotype, associated with, gene/protein),
-(effect/phenotype, parent-child, effect/phenotype),
-(effect/phenotype, phenotype absent, disease),
-(effect/phenotype, phenotype present, disease),
-(effect/phenotype, side effect, drug),
-(exposure, interacts with, biological_process),
-(exposure, interacts with, cellular_component),
-(exposure, interacts with, gene/protein),
-(exposure, interacts with, molecular_function),
-(exposure, linked to, disease),
-(exposure, parent-child, exposure),
-(gene/protein, associated with, disease),
-(gene/protein, associated with, effect/phenotype),
-(gene/protein, carrier, drug),
-(gene/protein, enzyme, drug),
-(gene/protein, expression absent, anatomy),
-(gene/protein, expression present, anatomy),
-(gene/protein, interacts with, biological_process),
-(gene/protein, interacts with, cellular_component),
-(gene/protein, interacts with, exposure),
-(gene/protein, interacts with, molecular_function),
-(gene/protein, interacts with, pathway),
-(gene/protein, ppi, gene/protein),
-(gene/protein, target, drug),
-(gene/protein, transporter, drug),
-(molecular_function, interacts with, exposure),
-(molecular_function, interacts with, gene/protein),
-(molecular_function, parent-child, molecular_function),
-(pathway, interacts with, gene/protein),
-(pathway, parent-child, pathway)]
 Relational query templates.
@@
 (anatomy → [expression present] → gene/protein → enzyme → drug):
-    "What is the drug that some genes or proteins act as an enzyme upon, where the genes or proteins are expressed in <anatomy>?",
+    "What is the drug that some genes or proteins act as an enzyme upon,
+     where the genes or proteins are expressed in <anatomy>?",
 (cellular_component → [interacts with] → gene/protein → [carrier] → drug):
@@
     "What anatomy does not express the genes or proteins associated with <effect/phenotype>?",
-(drug → [indication] → disease → [indication] → drug) & (drug → [synergistic interaction] → drug):
+(drug → [indication] → disease → [indication] → drug)
+ & (drug → [synergistic interaction] → drug):
     "Find drugs that has a synergistic interaction with <drug> and both are indicated for the same disease.",
-(pathway → [interacts with] → gene/protein → [interacts with] → pathway) & (pathway → [parent-child] → pathway):
+(pathway → [interacts with] → gene/protein → [interacts with] → pathway)
+    & (pathway → [parent-child] → pathway):
     "Find pathway that is related with <pathway> and both can [interacts with] the same gene/protein.",
-(gene/protein → [associated with] → disease → [associated with] → gene/protein) & (gene/protein → [ppi] → gene/protein):
+(gene/protein → [associated with] → disease → [associated with] → gene/protein)
+    & (gene/protein → [ppi] → gene/protein):
     "Find gene/protein that can interect with <gene/protein> and both are associated with the same disease.",
-(gene/protein → [associated with] → effect/phenotype → [associated with] → gene/protein) & (gene/protein → [ppi] → gene/protein):
+(gene/protein → [associated with] → effect/phenotype → [associated with] → gene/protein)
+    & (gene/protein → [ppi] → gene/protein):
     "Find gene/protein that can interect with <gene/protein> and both are associated with the same effect/phenotype."
 where []delimited-[][$\cdot$][ ⋅ ] denotes the relation type.
-Appendix B Prompts
-B.1 Extracting textual requirements
+Appendix C Prompts and LLM versions for Query Synthesization
+We summarize the LLM versions in Table 11.
+We chose these models based on their cost, how accurate they are, and whether they were the latest model during different parts of the project.
+While we used different LLMs, we checked each step separately to make sure the good quality in our benchmark datasets.
+Table 11: Summary of LLM Versions for Query Synthesization
+Step STaRK-Amazon STaRK-MAG STaRK-Prime
+Step 2: Extracting textual requirements gpt-3.5-turbo-16k claude-2.0 claude-2.0
+Step 3: Combining relational and textual requirements claude-2.0, gpt-4-0125-preview
+Step 4: Filtering additional answers claude-2.1, claude-2.0, claude-instant-1.2
+C.1 Extracting textual requirements
 {mdframed}
@@
 - description: <product description>
-- features:
-#1: <feature #1>
+- features: #1: <feature #1> ...
 - reviews:
@@
 text: <full review text>
-#2:
+#2: ...
 - Q&A:
@@
 answer: <answer to product-related question>
-#2:
+#2: ...
 Based on the given product information, you need to (1) identify the products generic category, (2) list all of the negative perspectives and their sources, and (2) extract up to five hard and five soft requirements relevant to customers interests along with their sources.
@@
 This is the information of the product that you need to write response for:
-Information:
 <product_doc>
@@
 Their questions tend to be straightforward, focusing on practical concerns on the symptons, effects, and inheritance etc., instead of the detailed mechanisms, which may also include more context.
-B.2 Combining relational and textual requirements
+C.2 Combining relational and textual requirements
 {mdframed}
@@
 Please output only the rewritten query:
-B.3 Filtering additional answers
+C.3 Filtering additional answers
 {mdframed}
@@
 Please provide only the JSON in your response. Response:"
+Appendix D Instructions for Generating Queries
+For the process of generating queries by human, the participants were given a list of entity IDs that we randomly sampled from the entire entity set.
+Then, they were asked to follow the following instructions with the support of our built interactive demo at https://huggingface.co/spaces/snap-stanford/SKB-Explorer.
+{mdframed}[backgroundcolor=gray!10, linecolor=black, linewidth=2pt, innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin=3pt, innerbottommargin=10pt] Task:
+1) Given the provided entity ID, review the associated document and any connected entities and multi-hop paths.
+2) Find interesting aspects of the entities by examining both their relational structures and the textual information available.
+3) Write your queries from these aspects such that the entity can satisfy all of them.
+Note:
+1) Please do not leak the name of the entity in the query.
+2) You can skip some queries if you think the knowledge involved is hard to understand.
+3) Feel free to be creative with content of your queries, you can also include additional context.
+There is NO restriction on how you express the queries.
+After that, we filtering the ground truth answers manually by human validation.

### v3

@@
 Answering real-world complex queries, such as complex product search, often requires accurate retrieval from semi-structured knowledge bases that involve blend of unstructured (e.g., textual descriptions of products) and structured (e.g., entity relations of products) information.
-However, previous works have mostly studied textual and relational retrieval tasks as separate topics.
+However, many previous works studied textual and relational retrieval tasks as separate topics.
 To address the gap, we develop STaRK, a large-scale Semi-structure retrieval benchmark on Textual and Relational Knowledge Bases.
-Our benchmark covers three domains/datasets: product search, academic paper search, and queries in precision medicine.
+Our benchmark covers three domains: product search, academic paper search, and queries in precision medicine.
 We design a novel pipeline to synthesize realistic user queries that integrate diverse relational information and complex textual properties, together with their ground-truth answers (items).
@@
 STaRK serves as a comprehensive testbed for evaluating the performance of retrieval systems driven by large language models (LLMs).
-Our experiments suggest that STaRK presents significant challenges to the current retrieval and LLM systems, indicating the demand for building more capable retrieval systems.
-The benchmark data and code are available on https://github.com/snap-stanford/stark.
+Our experiments suggest that STaRK presents significant challenges to the current retrieval and LLM systems, highlighting the need for more capable semi-structured retrieval systems.
 Figure 1: STaRK features queries on Semi-structured Knowledge Base (SKB) with textual and relational knowledge, with node entities as ground-truth answers.
@@
 bind interleukin-31 to activate the PI3K/AKT and MAPK pathways?
-Table 1: Example queries from STaRK, which involve semi-structured information, where the relational and textual aspects are highlighted.
+Table 1: STaRK QA examples which involve semi-structured (relational and textual) information.
 1 Introduction
@@
 Answering such queries is crucial for enhancing user experience, supporting informed decision-making, and preventing hallucination.
-To answer such queries the underlying knowledge can be represented in semi-structured knowledge bases (SKBs) [31, 43, 34], which integrate unstructured data, such as textual descriptions and natural language expressions (e.g., description of the tricycle), with structured data, like entity interactions on knowledge graph (e.g., a tricycle “brand” is Radio Flyer).
+To answer such queries, the underlying knowledge can be represented in semi-structured knowledge bases (SKBs) [36, 51, 41], which integrate unstructured data, such as natural language descriptions and expressions (e.g., description of the tricycle), with structured data, like entity interactions on knowledge graphs (e.g., a tricycle “brand” is Radio Flyer).
 This allows the SKBs to represent comprehensive knowledge in specific applications, making them indispensable in domains such as e-commerce [16], social media [32], and precision medicine [9, 19, 24].
 Limitations of prior works.
-Prior works and benchmarks have primarily focused on either purely textual queries on unstructured knowledge [18, 12, 23, 26, 46] or structured SQL or knowledge graph queries [2, 50, 48, 38], which are inadequate to study the complexities of retrieval tasks involving SKBs.
-Another line of works [10, 22, 48, 44, 45, 39] have explored the middle ground via multi-hop reasoning within single or across multiple documents.
-However, they are either limited in the span of knowledge or lack of diverse textual properties and explicit relational information.
+Prior works focused on either purely textual queries on unstructured knowledge [21, 15, 26, 30, 54, 13, 25, 56] or structured SQL [61, 60, 61, 60] or knowledge graph queries [5, 58, 46, 47, 59, 48, 8, 14, 17, 3], which are limited in the span of knowledge and inadequate to study the complexities of retrieval on SKBs.
 Recently, large language models (LLMs) have demonstrated significant potential on information retrieval tasks [62, 15, 31, 44].
-Nevertheless, it remains an open question how effectively LLMs can be applied to the specific challenge of retrieval from SKBs.
-It is also important to note that existing works focus mainly on general knowledge, e.g., from Wikipedia.
-In fact, the knowledge may commonly come from private sources, requiring information retrieval systems to operate on private SKBs.
-Therefore, there is a gap in our understanding of how current LLM retrieval systems handle the complex interplay between textual and relational requirements in queries that optionally involve private knowledge.
+Nevertheless, it remains an open question of how effectively LLMs can be applied to the challenging retrieval tasks on SKBs.
+Moreover, the existing works mainly focus mainly on general knowledge, e.g., from Wikipedia.
+However, the knowledge may commonly come from private sources, requiring retrieval systems to operate on private SKBs.
+Therefore, there is a gap of how current LLM retrieval systems handle the complex textual and relational requirements in queries that can involve private knowledge.
 Present work.
@@
 The key technical challenge that we solve is how to accurately simulate user queries on SKBs.
-This difficulty arises from the interdependence of textual and relational information, which can lead to challenges in precisely filtering through millions of candidates to construct the ground-truth answers.
-Additionally, ensuring that queries are relevant to real-world applications and resembles real-world scenarios adds further complexity to the benchmarking process.
+This difficulty arises from the interdependence of textual and relational information, which leads to challenges in precisely construct the ground-truth answers from millions of candidates.
+Additionally, ensuring that queries are useful and resembles real-world scenarios adds further complexity to the benchmarking process.
 We develop a novel pipeline that simulates user queries and constructs precise ground truth answers using three SKBs built from extensive texts and millions of entity relations from public sources.
 We validate the quality of queries in our benchmark through detailed analysis and human evaluation, focusing on their naturalness, diversity, and practicality.
-Furthermore, we incorporate 274 human-generated queries into our benchmark to compare with synthesized queries and enrich the testing scenarios.
+Furthermore, we incorporate 274 human-generated queries to compare with synthesized queries and enrich the testing scenarios.
 With STaRK, we delve deeper into retrieval tasks on SKBs, evaluate the capability of current retrieval systems, and provide insights for future advancement.
 Key features of STaRK are:
-  • Natural-sounding queries on SKBs (Table 1): Queries in our benchmark are crafted to incorporate rich relational information and complex textual properties.
-Additionally, these queries closely mirror the types of questions users would naturally ask in real-life scenarios, e.g., with flexible query formats and possibly with extra contexts.
+  • Natural-sounding queries on SKBs (Table 1): Queries in our benchmark incorporate rich relational information and complex textual properties.
+Additionally, these queries closely mirror the types of questions users would naturally ask in real-life scenarios, e.g., with flexible query formats and possibly with additional contexts.
   • Context-specific reasoning: The queries entail reasoning capabilities specific to the context.
@@
 Such reasoning capabilities are crucial for accurately interpreting and responding to the nuanced requirements of each query.
-  • Diverse domains: Our benchmark spans a wide range of domains, reflecting the diverse nature of real-world information needs.
-It covers three different knowledge bases1footnotetext: 1 We provide an interactive interface for exploring the SKBs at https://huggingface.co/spaces/snap-stanford/SKB-Explorer with applications on product recommendation, academic paper search, and precision medicine inquiry.
-Therefore, STaRK provides a comprehensive evaluation of retrieval systems across various contexts and knowledge domains.
-We also perform extensive experiments on each benchmark dataset for the prevailing retrieval systems.
-We highlight research challenges especially on the capability to handle textual and relational requirements and the latency of retrieval systems on large-scale SKBs that involve million-scale entities or relations.
-Finally, we provide insights into future directions to build more capable retrieval systems over private knowledge bases.
+  • Diverse domains: Our benchmark spans three knowledge bases***Explore the SKBs at https://stark.stanford.edu/skb_explorer.html for applications including product recommendation, academic paper search, and precision medicine inquiries.
+STaRK provides a comprehensive evaluation of retrieval systems across diverse contexts and domains.
+We conduct extensive experiments on LLM retrieval systems, highlighting challenges in handling textual and relational data and latency on large-scale SKBs with millions of entities or relations.
+Finally, we offer insights into building more capable retrieval systems to handle real-world complexity.
 2 Benchmarking Retrieval Tasks over Textual and Relational Knowledge
 Figure 2: Demonstration of Semi-structured Knowledge Bases, where each knowledge base combines both textual and relational information in a complex way, making the retrieval tasks challenging.
-Table 2: Data statistics of our constructed semi-structured knowledge bases
-#entity #relation avg. #entities #relations #tokens
-types types degree
-STaRK-Amazon 4 4 18.2 1,035,542 9,443,802 592,067,882
-STaRK-MAG 4 4 43.5 1,872,968 39,802,116 212,602,571
-STaRK-Prime 10 18 125.2 129,375 8,100,498 31,844,769
 2.1 Problem Definition
@@
 $D=\bigcup_{i\in V}D_{i}$ be the collection of free-form text documents associated with the nodes, where $D_{i}$ is the set of documents associated with node $i$.
-For example, the product knowledge graph in e-commerce can capture relationships between products and brands/colors/categories, and the corresponding text documents include product descriptions, reviews, etc..
+For example, the product knowledge graph in e-commerce can capture relationships between products and brands/colors/categories, and the corresponding text documents include product descriptions, reviews, etc.
 We define the tasks on our benchmark datasets as follows: Given the knowledge graph $G=(V,E)$, a collection of free text documents $D$, and a query $Q$, the output is a set of nodes $A\subseteq V$ such that for each node $i\in A$, it satisfies the relational requirements imposed by the structure of $G$ as specified in $Q$, and the associated documents $D_{i}$ satisfy the textual requirements specified in $Q$.
 2.2 Semi-structured Knowledge Bases (SKBs)
-As shown in Figure 2, we construct three SKBs from the relational structure between entities and the textual information associated with a subset of the entities.
-We present the statistics of the relational structure in Table 2 and introduce each SKB as follows:
+Table 2: Data statistics of our constructed semi-structured knowledge bases
+#entity #relation avg. #entities #relations #tokens
+types types degree
+STaRK-Amazon 4 5 18.2 1,035,542 9,443,802 592,067,882
+STaRK-MAG 4 4 43.5 1,872,968 39,802,116 212,602,571
+STaRK-Prime 10 18 125.2 129,375 8,100,498 31,844,769
+As shown Figure 2, we construct three large-scale SKBs with the relational and textual information with each entity. See Table 2 for the basic data statistics and Appendix A.1 for details.
 Amazon Semi-structured Knowledge Base.
-This knowledge base is derived from the Sports and Outdoors category from Amazon Product Reviews [13] and Amazon Question and Answer Data [29].
-The knowledge base features four entity types: product, brand, color, and category, and four types of relational information: also_bought, also_viewed between product entities, has_brand, has_color, has_category relations between product and brand, color, and category entities, respectively.
-In total, it comprises around 1.0M entities (product entities: 0.9M, brand entities: 0.1M, category entities: 1.4k, color entities: 1.7k) and 9.4M relations (also_bought: 2.8M, also_viewed: 1.9M, has_brand: 1.7M, has_category: 2.3M, has_color: 0.6M).
-We obtain the textual information by combining the meta data from Amazon Product Reviews with the customer Q&A records from Amazon Question and Answer Data.
-This provides a rich amount of textual data, including product titles, descriptions, prices, customer reviews, and related customer Q&A for each product.
-For other entities, we extract their names or titles as the textual attribute.
-Amazon SKB features an extensive textual data, which is largely contributed from customer reviews and Q&A.
+The SKB features four entity types: product, brand, color, and category, and five relation types: also_bought, also_viewed between product entities, and has_brand/color/category associated with the products.
+We derive the textual information of product nodes by combining Amazon Product Reviews [16] with Amazon Q&A Data [33].
+This provides a rich amount of texts, including product descriptions and customer reviews.
+For other entities, we extract their names or titles as the textual attributes.
+Amazon SKB features an extensive textual data largely contributed from customer reviews and Q&A.
 MAG Semi-structured Knowledge Base.
-This knowledge base is constructed from obgn-MAG [17], obgn-papers100M [17], and Microsoft Academic Graph (version 2019-03-22) [43, 37].
-The knowledge base contains around 1.9M entities under four entity types (author: 1.1M, paper: 0.7M, institution: 8.7K, field_of_study: 59.5k) and 39.8M relations under four relation types (author_writes_paper: 13.5M, paper_has_field_of_study: 14.5M, paper_cites_paper: 9.7M, author_affiliated_with_institution: 2.0M).
-We construct the knowledge base by selecting papers1 shared between obgn-MAG and obgn-papers100M.
-The relational structure comes from the extracted subgraph in obgn-mag based on the selected papers.
-The textual information such as titles and abstracts is sourced from obgn-papers100M.
-Additionally, we augment the textual data by integrating details from the Microsoft Academic Graph database, providing extra information like paper venue, author and institution names.
-This SKB demonstrates a large number of entities and relations associate with paper nodes, especially on citation and authorship relation types.
+This SKB includes node entities of paper, author, institute, and field_of_study.
+We derive its relational structure by extracting a subgraph from obgn-mag [20], which contains shared paper nodes with obgn-papers100M [20] and all non-paper nodes.
+We filter out non-English language papers as we only consider single-lingual queries.
+The paper documents include their titles and abstracts.
+Additionally, we integrating details from the Microsoft Academic Graph database (version 2019-03-22) [51, 45], providing extra textual information like paper venue, author and institution names.
+This SKB demonstrates a large number of relations associate with paper nodes, especially on citation and authorship relations.
 Prime Semi-structured Knowledge Base.
-We leverage the existing knowledge graph PrimeKG [6] which contains ten entity types including disease, drug, gene/protein, and eighteen relation types, such as associated_with, synergistic_interaction, indication, expression_present.
-The entity count in our knowledge base is approximately 129.3K, with around 8.1M relations.
-The details on entity numbers and relation tuples are available in Appendix B.3.
+We leverage the exisiting knowledge graph PrimeKG [9] which contains ten entity types including disease, gene/protein, and eighteen relation types, such as associated_with, indication.
 Compared to the Amazon and MAG SKBs, Prime SKB is denser and features a greater variety of relation types.
-While PrimeKG already provides text information on disease and drug entities, including drug mechanisms and disease descriptions, we additionally integrate the textual details from multiple databases for gene/protein and pathway entities such as genomic position, gene activity summary and pathway orthologous event.
-We include the public sources for SKBs construction in Table 10 (Appendix B).
-footnotetext: 1 We filter out non-English language papers as we only considers single-lingual queries.
+While PrimeKG provides text information on disease and drug entities, we integrate the data from multiple databases for gene/protein and pathway entities such as genomic position, gene activity summary and pathway orthologous event.
 2.3 Retrieval Tasks on Semi-structured Knowledge Bases
+Our retrieval benchmark (Table 3) consists of three novel retrieval-based question-answering datasets, each comprising synthesized train/val/test sets with $9$k to $14$k queries in total and a high-quality human-generate query set.
+The queries synthesize relational and textual knowledge, mirroring real-world queries in terms of natural-sounding property and flexible formats.
+STaRK-Amazon.
+The task aims at product recommendation, with a notable 68% of the synthesized queries yielding more than one ground truth answer.
+The dataset prioritizes customer-oriented criteria, highlighting textual elements such as product quality, functionality, and style.
+Moreover, it incorporate single-hop relational aspects (Appendix A.2) into the queries, including brand, category, and product connections (e.g., complementary or substitute items).
+The queries are framed in conversation-like formats, enriching the context and enhancing the dataset’s relevance to real-world scenarios.
+STaRK-MAG.
+Beyond the single-hop relational requirements in STaRK-Amazon, STaRK-MAG emphasizes the fusion between the textual requirements with multi-hop queries for precise academic paper search.
+For example, “Are there any papers from King’s College London” highlights the metapath (institution $\rightarrow$ author $\rightarrow$ paper) on the relational structure.
+We designed three single-hop and four multi-hop relational query templates (Appendix A.3).
+The textual aspects focus on the paper’s topic, methodology, and contribution etc.
+STaRK-Prime.
+The task is to answer complex biomedicine inquiries.
+For synthesized queries, we developed 28 multi-hop query templates (Appendix A.4) to cover various relation types and ensure their practical relevance.
+For example, the template “What is the drug that targets genes or proteins in <anatomy>?” aids precision medicine by identifying treatments targeted to specific anatomical areas.
+For drug, disease, gene/protein, and pathway entities, the queries are a hybrid of relational and textual requirements.
+For entities such as effect/phenotype, the queries rely solely on relational data due to limited textual information.
+We exhibit three distinct user roles – medical scientist, doctor, and patient – for generating queries about drug and disease, which diversify the language to comprehensively evaluate the retrieval systems.
 Table 3: Statistics on the STaRK benchmark datasets.
@@
 STaRK-MAG 84 34 3.26
-STaRK-Prime 109 43 2.61
-We develop three novel retrieval-based question-answering datasets.
-To highlight, the queries feature the synthesization between relational and textual knowledge.
-Moreover, to elevate their applicability in practical scenarios, these queries are designed to mimic real-life query patterns in terms of the natural-sounding property and flexible formats.
-We detail our retrieval benchmarks’ scale in Table 3 and introduce the specifics of each retrieval dataset as follows:
-STaRK-Amazon.
-This dataset comprises synthesized 9,100 customer-focused queries along with 81 human-generated queries for additional evaluation.
-The task aims at product searches, with a notable 68% of the synthesized queries yielding multiple answers.
-The dataset prioritizes customer-oriented criteria, highlighting textual elements such as product quality, functionality, and style.
-Additionally, it accentuates relational aspects, including brand, category, color and product connections (e.g., complementary or substitute items).
-The queries are framed in detailed, conversation-like formats, enriching the context and enhancing the dataset’s relevance to real-world scenarios.
-STaRK-MAG.
-This dataset comprises 13,323 synthesized and 84 human-generated queries targeting at academic paper search.
-Beyond the single-hop relational requirements on STaRK-Amazon, STaRK-MAG also emphasizes the fusion between the textual requirements with multi-hop queries.
-For example, “Are there any papers from King’s College London” highlights the metapath (institution $\rightarrow$ author $\rightarrow$ paper) on the relational structure.
-We designed three single-hop and four multi-hop relational query templates, cf.
-Appendix B.2.
-The textual aspects of the queries diversifies, focusing on different elements such as the paper’s topic, methodology, or the advancements it introduces, sourced mainly from the abstracts.
-STaRK-Prime.
-This dataset, comprising 11,204 synthesized queries and 109 human-generated queries across all ten entity types, incorporates single-hop and multi-hop relational information similar to STaRK-MAG.
-For synthesized queries, we developed 28 multi-hop query templates, detailed in Appendix B.3, to cover various relation types and ensure their practical relevance.
-For instance, the query “What is the drug that targets the genes or proteins expressed in <anatomy>?” serves applications in precision medicine and pharmacogenomics, aiding researchers and healthcare professionals in identifying drugs that act on genes or proteins associated with specific anatomical areas and enabling more targeted treatments.
-For entities like drug, disease, gene/protein, and pathway, the queries are a hybrid of relational requirements and textual requirements drawn from documents.
-A notable feature is the deliberate simulation of three distinct roles – medical scientist, doctor, and patient – for certain queries related to drugs and diseases.
-This is intended to diversify the language used across various user types and to evaluate the robustness of the retrieval system under varied linguistic scenarios.
-For entities such as effect/phenotype, the queries rely solely on relational data due to limited textual information in the knowledge base.
-We divide the synthesized queries on each dataset into training, validation, and testing subsets, with the ratios detailed in Table 3, while using the human-generated queries as additional sets for evaluation.
-For the STaRK-Amazon dataset, we randomly allocate a subset of queries with 20 or fewer answers to the validation and testing sets, while the remaining queries are assigned to the training set.
-For the STaRK-MAG and STaRK-Prime, we use random splits to divide the data.
+STaRK-Prime 98 41 2.77
+Figure 3: The construct pipeline to generate our semi-structured retrieval datasets.
 2.4 Benchmark Construction: Synthesized Queries
-Figure 3: The process of constructing semi-structured retrieval datasets involves four main steps: 1) Sample Relational Requirement: Based on relational templates, sample a relational requirement on a SKB.
-2) Extract Textual Properties: From a node that meets the relational requirement, extract relevant textual properties.
-3) Combine Information: Merge the relational information and textual properties to form a natural-sounding query.
-4) Construct Ground Truth Nodes: Check if nodes satisfy the textual properties using multiple language models to establish ground truth nodes.
-We present the novel pipeline that synthesizes the large-scale retrieval datasets.
-The core idea is to entangle relational information and textual properties and accurately construct ground truth answers.
-The construction of the retrieval datasets (Figure 3) generally involves four steps, and the specific processes vary depending on the characteristics of each dataset.
-These steps are as follows:
-  • 1) Sample Relational Requirements: Initially, we sample a relation template, such as “(a product) belongs to <brand>” and ground it into a relational requirement, e.g., “belongs to Radio Flyer” in Figure 3.
-Each relational requirement yields a set of candidate entities that meet the requirement, i.e., products belonging to Radio Flyer.
-The relation templates for the three SKBs are constructed with expert/domain knowledge, which are available in Appendix B.
-Each relation template in the SKBs serves a distinct and practical purpose.
-For instance, the template “What is the drug that targets the genes or proteins which are expressed in <anatomy>?” is particularly valuable for precision medicine inquiries.
-It aids in pinpointing medications aimed at specific genetic or protein targets within defined anatomical areas, thereby contributing to the development of targeted treatments.
-  • 2) Extracting Textual Properties: The relational requirement yields a set of satisfying candidates.
-Subsequently, we select one of the candidate nodes, referred to as the gold answer, and use LLMs to extract textual properties of interest from the its document.
-We extract the textual properties that align with the interests of specific roles, such as customers, researchers, or medical scientists, according to the specific SKB.
-For example, in Figure 3, we extract the phrases of functionality and user experience from a Radio Flyer product from different locations of the product document.
-  • 3) Combining Textual and Relational Information: After obtaining the textual and relational requirements, we synthesize them into a query using LLMs.
-We conduct two-stage synthesization using two different LLMs to avoid bias that might arise from relying on a single LLM and ensure a more diverse set of simulated queries.
-Specifically, this first-stage synthesization is guided by various criteria, such as natural sounding, adhering to the style of ArXiv searches, or aligning with specific roles.
-Furthermore, the second stage instructs the LLM to enrich the context and rephrase the language, demanding reasoning abilities in comprehending the requirements of the query.
-  • 4) Filtering Additional Answers: Finally, we employ multiple LLMs to verify if each remaining candidate, excluding the gold answer used for extracting textual properties, meets the additional textual requirement.
-Only candidates that pass the verification across all LLMs are included in the final ground truth answer set.
-This stringent verification ensures the accuracy of the nodes in the ground truth answers.
-To validate the precision of this filtering process, we compute the average precision of each gold answer passes the verification.
-The precision rates on STaRK-Amazon, STaRK-MAG, and STaRK-Prime are 86.6%, 98.9%, and 92.3%, respectively, demonstrating the effectiveness of our filtering approach in maintaining high-quality ground truth answers.
-This dataset construction pipeline is automatic, efficient, and broadly applicable to the SKBs in our formulation. We include all of the prompts and the LLMs versions in the above steps in Appendix C.
-2.5 Synthesized Data Distribution Analysis and Human Evaluation
+In Figure 3, we present a novel pipeline that synthesizes the SKB queries and automatically generates the ground truth answers.
+The key idea is to entangle relational and textual information during synthesis and disentangle them during answer filtering.
+It involves four steps as follows:
+  • 1) Sample Relational Requirements: For each query, we sample a practical relation template constructed with expert/domain knowledge, e.g., (a product) belongs to <brand>” and ground it with sampled entities (i.e., a specific brand), e.g., “belongs to Radio Flyer”.
+This relational requirement yields a set of candidate entities, i.e., products belonging to Radio Flyer.
+  • 2) Extracting Textual Properties: We randomly sample a candidate entity from the first step, referred to as the gold answer, from which LLMs extract properties that align with the interests of specific roles (e.g., customers, researchers, or doctors) in its textual document.
+In Figure 3, we extract multiple properties about the functionality and user experience from a Radio Flyer product.
+  • 3) Combining Textual and Relational Information: We use two LLMs to synthesize queries from textual properties and relational requirements, enhancing diversity and reducing bias arise from relying on a single LLM.
+The first LLM focuses on generating natural, role-specific, and style-consistent (e.g., ArXiv searches) queries.
+The second LLM enriches the context and rephrases queries, which poses the need for advanced reasoning to comprehend them under complex contexts.
+  • 4) Filtering Additional Answers: Finally, we employ multiple LLMs to verify if the candidates from the first step meet the extracted textual properties.
+Only candidates passing all LLM verifications are included in the final ground truth set.
+To assess the precision of this filtering mechanism, we compute the average ratios for the gold answers to be verified, which are 86.6%, 98.9%, and 92.3% on the three datasets, highlighting our efficacy in yielding high-quality ground truth answers.
+This dataset construction pipeline is automatic, efficient, and broadly applicable to the SKBs in our formulation. We include all of the prompts and the LLMs versions in the above steps in Appendix E.
 Figure 4: Distribution of query and answer lengths on STaRK datasets.
-Table 4: Query diversity measurement on STaRK. See Appendix A for the metric definition.
-Shannon Type-Token
-Entropy Ratio
+Table 4: Query diversity measurement on STaRK. See Appendix B for the metric definition.
+Shannon Entropy Type-Token Ratio
 STaRK-Amazon 10.39 0.179
@@
 STaRK-Prime 9.63 0.143
-Reference1 10.44 0.261
+Reference article 10.44 0.261
 Figure 5: Average relative composition of relational vs. textual information.
-We study the data distribution to understand the characteristics of our benchmark datasets and conduct human evaluation on the benchmark quality. We study the data distribution in three dimensions:
+2.5 Synthesized Data Distribution Analysis and Human Evaluation
   • Query and Answer Length.
-We analyze the query length distribution (measured in the number of words) and the number of ground truth answers for each query.
-The query length reflects the amount of context information provided by users, while the number of ground truth answers indicates the level of inclusiveness or ambiguity of queries within the specific SKB.
-As illustrated in Figure 4, all three datasets exhibit similar query length distributions, with most queries containing around 16 words.
-Queries with up to 50 words typically involve mentions of other entities, such as product or paper titles, or provide more detailed context, such as specific symptoms descriptions.
-Interestingly, the answer length distribution for STaRK-Amazon shows a more significant long-tail pattern, with approximately 22% of answers exceeding 30 entities, whereas the answer lengths for STaRK-Prime and STaRK-MAG are all within 20 entities.
-On average, as shown in Table 3, STaRK-Amazon presents an average answer length of 5.32, indicative of the e-commerce recommendation domain where general customer inquiries frequently result in diverse recommendations.
-While STaRK-Prime has the smallest average answer length, partially attributed to the smaller size of the Prime SKB compared to the other two SKBs.
+Query length (in words) reflects the amount of user-provided context information, while the number of answers indicates query ambiguity/concreteness.
+Figure 4 shows similar query length distributions across the datasets, with most queries around 16 words.
+Longer queries (up to 50 words) often mention other entities or provide detailed context.
+Notably, STaRK-Amazon has a significant long-tail pattern, with about 22% of the answers have more than 30 entities, reflecting diverse e-commerce recommendations and ambiguous user queries.
   • Query Diversity.
-A diverse set of queries pose challenges for broader applicability to meet varying user demands.
-Specifically, we measure query diversity using Shannon Entropy, which quantifies the uncertainty in the word distribution across all queries, and the Type-Token Ratio (TTR), which calculates the proportion of unique words to the total number of words.
-Higher values of Shannon Entropy and TTR indicate greater lexical diversity, reflecting a wider range of topics and linguistic expressions in the query set.
-As shown in Table 5, we observe high values of Shannon Entropy on all of the datasets.
-For reference, we compute the metrics for the wikipedia page of Barack Obama1footnotetext: 1https://en.wikipedia.org/wiki/Barack_Obama.
-While TTR is more sensitive to the document length, the TTR values consistently demonstrate a steady presence of unique words relative to the large length in our query document.
-  • Ratio of Relational vs.Textual Information.
-A key feature of our benchmark dataset is the composition of textual and relational information.
-Therefore, it is crucial to understand the proportionality between these two types of information in the final queries.
-We calculate the ratio of relational to textual requirements for each query by comparing the word count of relational information to the word count of extracted textual properties, both of which are meta data in the constructing process.
-We then average these ratios across each dataset.
-    Intuitively, the ratio serves as an approximation of the relative distribution based on the length of requirements, and does not directly reflect the importance of each type of information in determining the final answers.
-As shown in Figure 5, the ratios vary across different datasets, highlighting a differing emphasis on textual versus relational information.
-This distribution variation presents additional challenges for retrieval systems, requiring them to adapt to the dataset characteristic and specific balance of information.
-Table 5: Positive/Non-negative rates (%) from human evaluation.
-Naturalness Diversity Practicality
-STaRK-Amazon 73.6 / 89.5 68.4 / 89.5 89.5 / 94.7
-STaRK-MAG 94.7 / 100 73.7 / 84.2 68.4 / 84.2
-STaRK-Prime 67.8 / 92.8 71.4 / 82.1 71.4 / 89.3
-Average 78.7 / 94.1 71.0 / 85.3 76.4 / 89.4
+A diverse set of queries poses challenges for broader applicability to meet varying user demands.
+We measure query diversity using Shannon Entropy for word distribution and Type-Token Ratio (TTR) for unique words.
+Higher values indicate greater lexical diversity.
+Table 5 shows high Shannon Entropy and steady TTR across all datasets.
+For reference, we compute these metrics for the Wikipedia page of Barack Obamahttps://en.wikipedia.org/wiki/Barack_Obama.
+  • Proportionality of Relational vs.Textual Information.
+Our benchmark queries feature the composition of textual and relational information.
+To understand the distribution of information types, we calculate the average ratio of relational to textual requirements by word count in the queries across each dataset.
+Note that the ratios do not directly reflect their importance in determining final answers.
+Figure 5 shows varying ratios, which highlights different emphases on textual versus relational requirements and challenges retrieval systems to adapt to different distributions.
 Human evaluation.
-Moreover, we qualitatively assess sampled queries from our benchmark to determine if they are natural-sounding (resembling natural conversation), diverse (covering a wide range of question structures and complexity levels), and practical (relevant and useful in real-life situations), with participation from 63 individuals.
-The evaluation results are converted from a 5-point Likert-like scale to a positive/tie/negative scale for reporting.
-We report the positive and non-negative rates in Table 5.
-On average across the three datasets, 94.1%, 85.3%, and 89.4% of participants rated neutral or above in terms of naturalness, diversity, and practicality, respectively.
-The results justify the quality of our benchmark and highlight its potential for diverse and realistic retrieval tasks.
+We qualitatively assess sampled queries from our benchmark for naturalness (resembling natural conversation), diversity (covering various question structures and complexities), and practicality (relevance to real-world situations) with 63 participants.
+Evaluation results, converted from a 5-point Likert-like scale to a positive/tie/negative scale, show positive and non-negative rates in Table 10 (Appendix D.1).
+On average, 94.1%, 85.3%, and 89.4% of participants rated the queries neutral or above in naturalness, diversity, and practicality, respectively.
+These results validate the quality of our benchmark and its potential for diverse and realistic retrieval tasks.
 2.6 Benchmark Construction: Human-Generated Queries
-To enhance the practical relevance of our benchmark dataset, we engaged 31 participants, 22 of whom are native English speakers, to generate a total of 274 queries across three SKBs.
-The participants were given a list of entity IDs, randomly selected from the entire entity set.
-They followed detailed instructions (see Appendix D) using our interactive platform to explore the SKB data.
-Subsequently, we manually filtered the ground truth answers, ensuring their accuracy through human verification.
-We then analyze the commonality and difference between synthesized and human-generated queries.
+To enhance our benchmark’s practical relevance, we engaged 31 participants (22 native English speakers) to generate 263 queries across three SKBs following the detailed instructions (Appendix C) along with our interactive platform.
+We manually verified and filtered the ground truth answers to ensure the answer correctness.
+Table 3 shows the statistics of the human-generated datasets.
+Finally, we analyzed the commonalities and differences between synthesized and human-generated queries.
 Table 5: Comparison of Human-generated and Synthesized Queries
@@
 Complex Context Help me.
-I am trying to diagnose a patient with persistent joint pain, and I suspect a condition where the bone is dying due to compromised blood supply, often linked to factors like steroid use, alcohol abuse, or underlying diseases - what’s the name of this sneaky bone-killing culprit?
-I’m experiencing joint pain accompanied by swelling, fever… I’m concerned about medications aggravating my fuzzy eyesight and potential blood clotting complications.
-Could you recommend treatments effective for my symptoms while minimizing these specific side effects?
+I am trying to diagnose a patient with persistent joint pain, and I suspect a condition where the bone is dying due to compromised blood supply, often linked to factors like steroid use, … what’s the name of this sneaky bone-killing culprit?
+I’m experiencing joint pain accompanied by swelling… I’m concerned about medications aggravating my fuzzy eyesight and potential blood clotting complications.
+Could you recommend treatments while minimizing these side effects?
 Commonality.
-We present the statistics of the human-generated datasets in Table 3.
-The synthesized and human-generated queries exhibit a similar level of ambiguity, resulting in a comparable number of answers.
-Moreover, we find that most styles and formats of human-generated queries are represented in our synthesized query dataset.
-In Table 6, we demonstrate three pairs of human-generated and synthesized queries that display similarities such as short and straightforward product queries, inquiries targeting specific authors and fields, and complex contextual queries in STaRK-Prime.
+The number of answers of synthesized and human-generated queries are comparable, indicating a similar level of query ambiguity.
+Moreover, we observe that most styles of human-generated queries are covered in the synthesized dataset.
+For example, Table 5 highlights their similarities in short product queries, specific author/field inquiries, and complex contextual queries.
 Difference.
-We find that human-generated queries often exhibit more unique expressions compared to synthesized queries, such as "Give me a fat cross and road tire that works with my Diamondback bicycle tube" and "this sneaky bone-killing culprit" (see the last row of Table 6).
+We find that human-generated queries often exhibit more unique expressions compared to synthesized ones, such as "Give me a fat cross and road tire that works with my Diamondback bicycle tube" and "this sneaky bone-killing culprit".
 This discovery suggests a future direction for our benchmark to incorporate modern and dynamic language nuances.
 3 Experiments
-In this section, we aim to explore the following questions:
-  • Q1: How well do current retrieval approaches perform on our benchmark datasets?
-  • Q2: What challenges are revealed by the experiments, and more importantly, what are the potential future directions for improvement?
-3.1 Retrieval Models
-Vector Similarity Search (VSS).
-VSS embeds both the query and the concatenated textual and relational information of each candidate entity.
-Then, the similarity score is computed based on cosine similarity between the query and candidate embeddings.
-We use the text-embedding-ada-002 model from OpenAI for generating the embeddings.
-Multi-Vector Similarity Search (Multi-VSS).
-Multi-VSS represents candidate entities with multiple vectors, capturing detailed features for complex retrieval tasks.
-Here, we chunk and separately embed the textual and relational information of each candidate entity.
-The final score is aggregated from the similarities between the chunks and the query.
-Dense Retriever.
-Dense Retriever finetunes a query encoder and a document encoder separately using the query answer pairs from the training dataset.
-We optimize the encoder using contrastive loss.
-Specifically, the positive pairs are constructed from the training questions and their ground truth answers, while we construct 20 hard negative answers for each query from the top false positive predictions from VSS.
-We use roberta-base as the base model and finetune the encoder over the entire training split for each dataset.
-QAGNN [49].
-QAGNN constructs a graph where nodes include entities found in the question or answer choices, incorporating their neighboring nodes as intermediate context.
-It extracts a subgraph from a larger knowledge graph, enabling a comprehensive understanding by leveraging the relational information from the knowledge graph.
-The approach integrates this with semantic embeddings from a language model, jointly modeling both relational and semantic information to enhance the question-answering modeling.
-VSS + LLM Reranker [9, 55].
-This method improves the precision of the top-$v$ results from VSS by reranking them using language models, taking advantage of their advanced language understanding capabilities.
-For this purpose, we employ two different language models: GPT-4-turbo (gpt-4-1106-preview) and Claude3 (claude-3-opus).
-We set $v=20$ for synthesize queries and $v=10$ for human-generated queries to balance between precision and computational efficiency.
-Specifically, we construct a prompt that instructs the language models to assign a score between 0 and 1 to a node, based on its combined textual and relational information, with certain criteria provided for rating the node.
-Due to the high retrieval cost, we randomly sample 10% of the testing queries on each dataset for evaluation.
-3.2 Evaluation Metrics
-We use the following metrics for a holistically evaluating the model performance.
-Hit@$k$.
-The Hit@$k$ metric assesses whether the correct item is among the top-$k$ results from the model.
+3.1 Baseline Retrieval Models and Evaluation Metrics
+We extensively evaluate five classes of retrieval models described below.
+  • Sparse Retriever: BM25 [40] is a traditional yet powerful sparse retrieval method based on term frequency-inverse document frequency (TF-IDF).
+It computes relevance scores by considering the frequency of query terms in documents, adjusted for term rarity and document length.
+  • Small Dense Retrievers: DPR [27], ANCE [53], and QAGNN [57].
+These compact models generate dense embeddings for both queries and documents, computing retrieval scores based on embedding similarities.
+They serve as baselines for comparison with LLM-based dense retrievers.
+  • LLM-based Dense Retrievers: text-embedding-ada-002 (abbrev.
+ada-002) [37], voyage-large-2-instruct (abbrev.
+voyage-l2-instruct) [2], LLM2Vec-Meta-Llama-3-8B-Instruct-mntp (abbrev.
+LLM2Vec) [4], and GritLM-7b [34].
+These models leverage LLMs to generate dense embeddings that are more contextually expressive.
+  • Multivector Retrievers: multi-ada-002 [37] and ColBERTv2 [42].
+Beyond ada-002 which represents a document as an embedding, multi-ada-002 splits each document into overlapping chunks and embeds them using the same encoder as the query.
+Similarity scores between the query and chunks are aggregated using the average of the top-3 similarities, which we found to perform best.
+ColBERTv2 represents each document as multiple token-level embeddings for fine-grained matching, capturing richer semantic information.
+  • LLM Rerankers: Claude3 and GPT-4 rerankers [12, 63].
+These models improve the precision of top-$k$ ada-002 results by reranking them using large language models.
+We employ GPT-4-turbo (gpt-4-1106-preview) and Claude3 (claude-3-opus), setting $k=20$ for synthesized queries and $k=10$ for human-generated queries.
+Given a query, the LLMs assign a satisfaction score from 0 to 1 to each candidate entity based on textual and relational information.
+Due to high computational costs, we evaluate these rerankers on a random 10% sample of test queries.
+The performance of these models are measured using standard retrieval metrics below.
+  • Hit@$k$ assesses whether the correct item is among the top-$k$ results from the model.
 We used $k=1$ and $k=5$ for evaluation.
 At $k=1$, it evaluates the accuracy of the top recommendation; at $k=5$, it examines the model’s precision in a wider recommendation set.
-It’s a straightforward yet effective way to measure the immediate relevance of answers.
-Recall@$k$.
   • Recall@$k$ measures the proportion of relevant items in the top-$k$ results.
 For synthesized queries, $k=20$ is used, as the answer length of all of the queries in our benchmarks are equal or smaller then $20$.
-For human-generated queries, Recall@20 is not reported as the number of responses in the human-generated subset may exceed 20.
 This metric offers insight into the model’s ability to identify all relevant items, particularly in scenarios where missing any could be critical.
-Mean Reciprocal Rank (MRR).
-MRR is a statistic for evaluating the average effectiveness of a predictive model.
+  • Mean Reciprocal Rank (MRR) is a statistic for evaluating the average effectiveness of a predictive model.
 It calculates the reciprocal of the rank at which the first relevant item appears in the list of predictions.
-MRR is particularly useful for understanding the model’s performance at presenting the correct item at an early stage, without the need for a predefined $k$.
 This metric emphasizes the importance of the rank of the first correct answer, which is crucial in many practical applications where the first correct answer is often the most impactful.
-3.3 Analysis
-Table 7: Testing results on STaRK-Syn(thesized). (*) indicates partial evaluation (10% of the testing queries) due to the high latency and cost of the methods.
-Dense Retriever QAGNN VSS Multi-VSS VSS+Claude3* VSS+GPT4*
-(roberta) (roberta) (ada-002) (ada-002) Reranker Reranker
-STaRK-Amazon Hit@1 0.1529 0.2656 0.3916 0.4007 0.4549 0.4479
-Hit@5 0.4793 0.5001 0.6273 0.6498 0.7113 0.7117
-Recall@20 0.4449 0.5205 0.5329 0.5512 0.5377 0.5535
-MRR 0.3020 0.3775 0.5035 0.5155 0.5591 0.5569
-STaRK-MAG Hit@1 0.1051 0.1288 0.2908 0.2592 0.3654 0.4090
-Hit@5 0.3523 0.3901 0.4961 0.5043 0.5317 0.5818
-Recall@20 0.4211 0.4697 0.4836 0.5080 0.4836 0.4860
-MRR 0.2134 0.2912 0.3862 0.3694 0.4415 0.4900
-STaRK-Prime Hit@1 0.0446 0.0885 0.1263 0.1510 0.1779 0.1828
-Hit@5 0.2185 0.2135 0.3149 0.3356 0.3690 0.3728
-Recall@20 0.3013 0.2963 0.3600 0.3805 0.3557 0.3405
-MRR 0.1238 0.1473 0.2141 0.2349 0.2627 0.2655
+3.2 Results and Analysis
+Table 6: Testing results on STaRK-Syn(thesized).
+STaRK-Amazon STaRK-MAG STaRK-Prime
+Hit@1 Hit@5 R@20 MRR Hit@1 Hit@5 R@20 MRR Hit@1 Hit@5 R@20 MRR
+Full Testing Dataset
+BM25 44.94 67.42 53.77 55.30 25.85 45.25 45.69 34.91 12.75 27.92 31.25 19.84
+DPR (roberta) 15.29 47.93 44.49 30.20 10.51 35.23 42.11 21.34 4.46 21.85 30.13 12.38
+ANCE (roberta) 30.96 51.06 41.95 40.66 21.96 36.50 35.32 29.14 6.53 15.67 16.52 11.05
+QAGNN (roberta) 26.56 50.01 52.05 37.75 12.88 39.01 46.97 29.12 8.85 21.35 29.63 14.73
+ada-002 39.16 62.73 53.29 50.35 29.08 49.61 48.36 38.62 12.63 31.49 36.00 21.41
+voyage-l2-instruct 40.93 64.37 54.28 51.60 30.06 50.58 50.49 39.66 10.85 30.23 37.83 19.99
+LLM2Vec 21.74 41.65 33.22 31.47 18.01 34.85 35.46 26.10 10.10 22.49 26.34 16.12
+GritLM-7b 42.08 66.87 56.52 53.46 37.90 56.74 46.40 47.25 15.57 33.42 39.09 24.11
+multi-ada-002 40.07 64.98 55.12 51.55 25.92 50.43 50.80 36.94 15.10 33.56 38.05 23.49
+ColBERTv2 46.10 66.02 53.44 55.51 31.18 46.42 43.94 38.39 11.75 23.85 25.04 17.39
+Random 10% Sample
+BM25 42.68 67.07 54.48 54.02 27.81 45.48 44.59 35.97 13.93 31.07 32.84 21.68
+DPR (roberta) 16.46 50.00 42.15 30.20 11.65 36.84 42.30 21.82 5.00 23.57 30.50 13.50
+ANCE (roberta) 30.09 49.27 41.91 39.30 22.89 37.26 44.16 30.00 6.78 16.15 17.07 11.42
+QAGNN (roberta) 25.00 48.17 51.65 36.87 12.03 37.97 47.98 28.70 7.14 17.14 32.95 16.27
+ada-002 39.02 64.02 49.30 50.32 28.20 52.63 49.25 38.55 15.36 31.07 37.88 23.50
+voyage-l2-instruct 43.29 67.68 56.04 54.20 34.59 50.75 50.75 42.90 12.14 31.42 37.34 21.23
+LLM2Vec 18.90 37.80 34.73 28.76 19.17 33.46 29.85 26.06 9.29 20.7 25.54 15.00
+GritLM-7b 43.29 71.34 56.14 55.07 38.35 58.64 46.38 48.25 16.79 34.29 41.11 24.99
+multi-ada-002 40.85 62.80 52.47 51.54 25.56 50.37 53.03 36.82 15.36 32.86 40.99 23.70
+ColBERTv2 44.31 65.24 51.00 55.07 31.58 47.36 45.72 38.98 15.00 26.07 27.78 19.98
+Claude3 Reranker 45.49 71.13 53.77 55.91 36.54 53.17 48.36 44.15 17.79 36.90 35.57 26.27
+GPT4 Reranker 44.79 71.17 55.35 55.69 40.90 58.18 48.60 49.00 18.28 37.28 34.05 26.55
 Table 7: Testing results on STaRK-Human(-Generated).
-VSS Multi-VSS VSS+Claude3 VSS+GPT4
-(ada-002) (ada-002) Reranker Reranker
-STaRK-Amazon Hit@1 0.3950 0.4691 0.5602 0.5432
-Hit@5 0.6419 0.7284 0.7284 0.7284
-MRR 0.5265 0.5874 0.6439 0.6274
-STaRK-MAG Hit@1 0.2857 0.2381 0.3810 0.3690
-Hit@5 0.4167 0.4167 0.4643 0.4523
-MRR 0.3581 0.3143 0.4243 0.4031
-STaRK-Prime Hit@1 0.2120 0.2567 0.2844 0.2844
-Hit@5 0.4037 0.4037 0.4771 0.4863
-MRR 0.2984 0.3377 0.3623 0.3617
+STaRK-Amazon STaRK-MAG STaRK-Prime
+Method Hit@1 Hit@5 R@20 MRR Hit@1 Hit@5 R@20 MRR Hit@1 Hit@5 R@20 MRR
+BM25 27.16 51.85 29.23 18.79 32.14 41.67 32.46 37.42 22.45 41.84 42.32 30.37
+DPR (roberta) 16.05 39.51 15.23 27.21 4.72 9.52 25.00 7.90 2.04 9.18 10.69 7.05
+ANCE (roberta) 25.93 54.32 23.69 37.12 25.00 30.95 27.24 27.98 7.14 13.27 11.72 10.07
+QAGNN (roberta) 22.22 49.38 21.54 31.33 20.24 26.19 28.76 25.53 6.12 13.27 17.62 9.39
+ada-002 39.50 64.19 35.46 52.65 28.57 41.67 35.95 35.81 17.35 34.69 41.09 26.35
+voyage-l2-instruct 35.80 62.96 33.01 47.84 22.62 36.90 32.44 29.68 16.33 32.65 39.01 24.33
+LLM2Vec 29.63 46.91 21.21 38.61 16.67 28.57 21.74 21.59 9.18 21.43 26.77 15.24
+GritLM-7b 40.74 71.60 36.30 53.21 34.52 44.04 34.57 38.72 25.51 41.84 48.10 34.28
+multi-ada-002 46.91 72.84 40.22 58.74 23.81 41.67 39.85 31.43 24.49 39.80 47.21 32.98
+ColBERTv2 33.33 55.56 29.03 43.77 33.33 36.90 30.50 35.97 15.31 26.53 25.56 19.67
+Claude3 Reranker 53.09 74.07 35.46 62.11 38.10 45.24 35.95 42.00 28.57 46.94 41.61 36.32
+GPT4 Reranker 50.62 75.31 35.46 61.06 36.90 46.43 35.95 40.65 28.57 44.90 41.61 34.82
 Results on synthesized queries.
-Table 7 shows the experiment results on the synthesized testing datasets.
-Firstly, we notice a significant performance gap between models trained or fine-tuned on our training datasets, such as Dense Retriever and QAGNN, and models like VSS and Multi-VSS that use embeddings generated from text-embedding-ada-002.
-For the dense retriever, training the encoders proves challenging when handling both textual and relational information as text, leading to training challenges.
-For QAGNN, it training challenges due to its high computational demands during training.
-Finally, the performance of VSS and Multi-VSS demonstrates the superior pre-trained embeddings obtained from a large corpus.
-Further, the Multi-VSS strategy, which breaks down the documents into chunks and embeds them into embeddings separately to capture more fine-grained context, generally outperforms VSS.
-However, it falls short in the Hit@1 metric on the STaRK-MAG dataset, likely due to this dataset’s property for understanding more extended contexts or potential interaction or reasoning between different chunks.
-Significantly, the methods combining VSS with an LLM reranker consistently outperform other approaches across most metrics.
-Notably, reranking the top 20 predictions with an LLM leads to significant improvements in Hit@$k$ and MRR metrics compared to using VSS alone.
-This improvement underscores the limitations of traditional embedding methods, which lack the advanced reasoning capabilities provided by LLMs.
-However, even for VSS+GPT4 Reranker, we observe insufficient performances, e.g., Hit@1 is only around 18% on STaRK-Prime and 41% on STaRK-MAG , meaning that for a significant portion of queries, the top-ranked answer is not the correct one.
-Moreover, Recall@20 metric values are lower than 60% across all of the datasets, which indicates the ranking results may not be comprehensive, potentially missing relevant answers in specific applications.
-Also, we notice that the MRR scores generally follow a similar trend as the Hit@1 metric.
-However, the MRR values remain relatively low, particularly for STaRK-Prime, indicating large space for further optimization in the ranking process.
+Table 6 presents the results on both the full synthesized test sets and random 10% samples from these sets.
+In both cases, BM25, despite its simplicity, proves to be a strong baseline, outperforming the dense retrieval models such as ANCE.
+We observe that finetuned DPR and QAGNN, exhibit insufficient performance.
+This underperformance is likely due to their relatively small model sizes and the risk of overfitting during training.
+These issues present challenges in effectively training the models on SKBs, where the entity documents can be hard to differentiate without capturing detailed information.
+Among the larger models, ada-002 benefits from superior pretrained embeddings and significantly outperforms LLM2Vec by a large margin.
+GritLM-7b delivers excellent performance, surpassing the ada-002 model overall.
+In contrast, LLM2Vec underperforms due to its limited context length, which is insufficient for encoding the lengthy documents in the SKBs.
+For multivector retrievers, we found that multi-ada-002 generally outperforms ada-002, indicating that using multiple vectors per document enhances retrieval effectiveness.
+Similarly, fine-grained representation allows ColBERTv2 to capture subtle semantic nuances between queries and documents, leading to largely improved retrieval accuracy.
+However, both GritLM-7b and ColBERTv2 generally underperform compared to the rerankers on the random split, especially in terms of Hit@k metrics.
+This suggests that while these dense retriever models effectively capture semantic information, they may not fully grasp the nuanced relevance judgments required for top-tier retrieval performance.
+The rerankers, utilizing powerful LLMs like GPT-4 (gpt-4-1106-preview) and Claude3 (claude-3-opus), excel by re-evaluating the top candidates and assigning satisfaction scores based on a deeper understanding of the query and document content.
+This process allows them to better discern subtle contextual cues and relational information that dense retrievers might overlook.
+Consequently, LLM rerankers enhance retrieval precision at the top ranks.
+Finally, regardless of the higher computational costs of the rerankers, their performance remains suboptimal.
+For instance, the Hit@1 scores for the GPT-4 reranker are only about 18% on STaRK-Primeand 41% on STaRK-MAG, indicating that the top-ranked answers are frequently incorrect.
+Similarly, the Recall@20 metrics are below 60% across all datasets, with the GPT-4 reranker achieving Recall@20 scores of 55% on STaRK-Amazon, 49% on STaRK-MAG, and 34% on STaRK-Prime.
+This suggests that the ranking results miss a significant portion of relevant answers.
+The MRR scores are also relatively low, especially for STaRK-Prime, where the GPT-4 reranker attains an MRR of only around 27%.
+The insufficient performances may be attributed to the complexity and diversity of queries in SKBs, where nuanced understanding and detailed contextual information are crucial.
+These findings highlight significant room for improvement in the ranking process.
 Results on human-generated dataset.
-We further evaluate the well-performing methods on human-generated queries.
-Table 8 presents the testing results on the human-generated dataset.
-Reranker methods continue to outperform other methods, confirming their superior reasoning and context understanding capabilities.
-Compared to synthesized datasets, the performance on human-generated queries shows highly correlated trends but generally higher metrics.
-This suggests that the synthesized datasets could be more difficult than the human-generated ones, offering challenging tasks and shedding light on interesting aspects of complex queries.
+Table 7 presents the testing results on the human-generated datasets.
+For example, the rerankers consistently outperform others, showing their reasoning and context understanding ability.
+Compared to the synthesized datasets, the performance on human-generated queries is generally higher for most models, but the overall trends remain consistent.
+This indicates that synthesized datasets may be more challenging, highlighting the complexity of the tasks on our synthesized queries.
+Another interesting observation is that the performance of the rerankers is particularly strong on human-generated queries, which may contain more nuanced language and diverse expressions.
+This suggests that rerankers excel in interpreting and leveraging the richness of human language to improve retrieval accuracy.
 Table 8: Latency (s) of the retrieval systems on STaRK.
-Dense Retriever QAGNN VSS Multi-VSS VSS+Claude VSS+GPT4
+DPR QAGNN ada-002 multi-ada-002 Claude3 Reranker GPT4 Reranker
 STaRK-Amazon 2.34 2.32 5.71 4.87 27.24 24.76
@@
 Average 1.40 1.65 2.83 2.97 26.33 25.05
+Figure 6: A case study on STaRK-MAG shows that ada-002 overranks non-ground truth papers $C$ and $D$ due to repeated keywords in the relational information “cites paper”.
+After reranking with Claude3, it correctly prioritizes ground truth papers $A$ and $B$ with accurate reasoning and analysis.
 Retrieval latency.
-Another aspect vital for the practical application of retrieval systems is their latency.
-As shown in Table 9, we test the latency of the retrieval methods on a single NVIDIA A100-SXM4-80GB.
-Specifically, the Dense Retriever and QAGNN models exhibit lower average latency, making them more suitable for time-sensitive applications.
-In contrast, the VSS and Multi-VSS models show moderate latency, but when combined with LLM rerankers, such as Claude and GPT-4, the latency significantly increases.
-Therefore, in the context of semi-structured retrieval tasks, the balance between enhanced accuracy and increased latency is crucial.
-Effective retrieval methods should consider such trade-offs, particularly when addressing complex queries involving reasoning over both textual and relational information.
-Figure 6: A case study on STaRK-MAG, where VSS mistakenly ranks non-ground truth papers C and D higher due to the repeated key words in the relational information "cites paper." After reranking with Claude, it correctly prioritizes the ground truth papers A and B.
-This correction is attributed to the more accurate reasoning and analysis of the combined textual and relational information.
+Latency is crucial for practical retrieval systems, as users expect quick responses.
+As shown in Table 8, we evaluated the latency of various models using a single NVIDIA A100-SXM4-80GB GPU.
+We observed that the DPR and QAGNN models exhibit lower average latency, making them suitable for time-sensitive applications.
+In contrast, the ada-002 and multi-ada-002 models have moderate latency due to multiple API calls.
+However, when combined with LLM rerankers, the latency increases significantly due to the computational demands of these large models.
+Therefore, it is important to balance accuracy and latency, especially for complex queries that require advanced reasoning capabilities.
 Case study.
-Finally, we highlight the importance of reasoning ability in achieving good retrieval results on our benchmark datasets.
-In Figure 6, we conduct a case study to compare the performance of VSS and VSS+Claude3 Reranker.
-Specifically, we examine a user query that requests papers from an institution on a particular topic.
-In this scenario, VSS fails to adequately address the relational requirement due to directly embedding the entire documents without detailed analysis.
-As a result, papers containing frequently repeated keywords, such as “nonlinear modeling” and “piezoelectric elements”, are mistakenly assigned high scores.
-However, the result improves significantly after the reranking with an LLM, where the LLM is equipped to reason the relationship between the query and the information contained in each paper, making the scores more accurately reflect the papers’ relevance to the query.
-This case study underscores the need for reasoning ability in grasping the complexities of queries.
-As embedding methods may capture some aspects of the query, they often fail to capture the detailed interaction between textual and relational information.
-Thus, integrating LLMs with advanced reasoning into the retrieval process can be crucial for the retrieval accuracy.
+To highlight the importance of reasoning ability for achieving good performance on our benchmark, we present a case study in Figure 6, comparing the ada-002 model with the Claude3 Reranker.
+In this example, the query requests papers from a specific institution on a particular topic.
+The ada-002 model fails to address the relational aspect of the query because it embeds entire documents without detailed analysis.
+This leads to high relevance scores for irrelevant papers that frequently mention keywords like "nonlinear modeling" and "piezoelectric elements" but do not satisfy the relational requirement.
+In contrast, the LLM reranker significantly improves the results by reasoning about the relationship between the query and each paper, resulting in scores that more accurately reflect relevance.
+This underscores the need for reasoning ability to grasp query complexities.
 4 Related Work
-We identify three primary categories on retrieval-oriented question answering (QA) datasets based on the knowledge sources they utilize: unstructured, structured, and semi-structured knowledge bases.
-We discuss the differences and limitations in the previous works that we seek to address.
 Unstructured QA Datasets.
-The line of work explores retrieving answers from a single document [33] or multiple document sources [42, 47, 22, 44, 10].
-For example, SQuAD [33] focuses on extracting answers from a single document, testing model’s ability to understand and retrieve information within a specific context.
-For multi-document retrieval, HotpotQA [47] demands reasoning over multiple documents and includes sentence-level supporting facts, enhancing the complexity of the retrieval task.
-TriviaQA [22] combines questions from trivia games with evidence documents, evaluating both retrieval and reasoning over multiple sources.
-Moreover, some works [30, 25] leverage generated results by search engines as ground truth answers or information sources for retrieval.
-However, unstructured QA datasets or knowledge bases often lack the depth of relational reasoning commonly required in answering complex user queries.
+This research domain consists of methods for retrieving answers from unstructured text, either from a single document [39] or multiple documents [50, 55, 25, 52, 13].
+For instance, SQuAD [39] is designed for answer extraction within a specific document, while approaches like HotpotQA [55] and TriviaQA [25] extend to multi-document contexts.
+Additionally, some studies utilize search engine outputs as a basis or supplementary data for question answering [35, 29].
+However, unstructured QA datasets often lack the depth of relational reasoning commonly required in answering complex user queries.
 In contrast, STaRK contains queries demanding multi-hop relational reasoning to challenge model’s ability of handling structured information.
 Structured QA Datasets.
-Within this category, the datasets challenge models to retrieve answers from structured knowledge bases, either knowledge graphs [3, 51, 40, 4, 5, 11, 14] or tabular data [53, 52].
-For example, ComplexWebQuestions [40] challenge models’ abilities to decompose complex queries and interpret queries using entities and relationships defined in knowledge graphs.
-GraphQA [14] textualizes graphs by describing node attributes and edges with natural language, offering language models a way to handle KBQA.
-For datasets with tabular data, WikiSQL [53] and Spider [52] both address the task of converting natural language questions to SQL queries, with WikiSQL concentrating on single-table databases and Spider extending to complex multi-tables.
-While these datasets primarily focus on relational information, the lack of textual information limits questions within predefined relationships and entities, which constrains the breadth of available information.
-STaRK extends this line of work by integrating textual information into knowledge bases and queries, resulting in more comprehensive and nuanced responses with extensive insights drawn from abundant textual data.
+These datasets challenge models to derive answers from structured sources such as knowledge graphs [6, 59, 48, 7, 8, 14, 17] or tabular data [61, 60].
+ComplexWebQuestions [48] and GraphQA [17] propose challenges in interpreting complex queries and textualizing graph structures in KBQA, respectively.
+For tabular data, WikiSQL [61] focuses on translating queries to SQL for single-table databases, whereas Spider [60] tackles multi-table scenarios.
+Despite the emphasis on relational data, the restriction to predefined entities and relationships limits the scope of queries.
+STaRK integrates textual content within structured frameworks to enhance the depth and breadth of information retrieval, promoting richer and more nuanced understanding from extensive textual data.
 Semi-Structured QA Datasets.
-Another line of work focuses on semi-structured QA with tabular and textual data.
-WikiTableQuestions [32] emphasizes comprehension of table structure and textual content, reflecting the semi-structured nature of many real-world datasets.
-TabFact [7] and HybridQA [8] combine textual and tabular data, requiring models to validate statements or answer questions using both data types.
-TabMCQ [20] introduces multiple-choice questions based on tables, adding complexity in understanding and reasoning over semi-structured data.
-However, different from this work, these existing datasets mainly focus on tables as structured sources, which may not incorporate rich relational information that naturally exists between entities.
-Additionally, previous attempts to combine text and tabular data through external web links or additional text columns often result in data structures that are hard to navigate.
-Therefore, STARK goes beyond these limitations by integrating both relational and textual data in a more cohesive manner, allowing for seamless navigation and retrieval across complex semi-structured knowledge bases.
-This integration supports richer relational reasoning and more comprehensive textual data handling, making STARK a more effective benchmark for evaluating the capabilities of retrieval systems in scenarios that closely mirror real-world information needs.
-5 Conclusion
-We introduce STaRK, the first benchmark designed to thoroughly evaluate the capability of retrieval systems driven by LLMs in handling semi-structured knowledge bases (SKBs).
-This benchmark features a diverse set of queries that are semi-structured and natural-sounding, requiring context-specific reasoning across various domains, thereby setting a new standard for assessing retrieval systems in the context of SKBs.
-We utilize public datasets to construct three SKBs and develop an automated, general pipeline to simulate user queries that mimic real-world scenarios.
-Moreover, we augment our datasets with human-generated queries and provide reference to the synthesized queries.
-Our extensive experiments on STaRK reveal significant challenges for current models in handling both textual and relational information effectively and flexibly.
-Overall, STaRK offers valuable opportunities for future research to advance the field of complex and multimodal retrieval systems, where reducing retrieval latency and incorporating strong reasoning ability into the retrieval process are identified as two prospective future directions.
+This category merges tabular and textual data, presenting challenges in semi-structured data comprehension.
+WikiTableQuestions [38] stresses the integration of table structures with textual elements.
+TabFact [10], HybridQA [11], and TabMCQ [23] extend this by combining validation of textual statements with tabular reasoning.
+However, datasets leveraging tables as structured frameworks often lack in depicting the rich relational dynamics among entities.
+Moreover, prior efforts to link textual and tabular information via external sources have led to cumbersome data constructs.
+Addressing these challenges, STaRK enhances integration, allowing for flexible navigation and advanced retrieval within complex semi-structured knowledge bases, and facilitating more effective relational reasoning and text handling.
+5 Conclusion and Future Work
+We introduce STaRK, the first benchmark to thoroughly evaluate LLM-driven retrieval systems for semi-structured knowledge bases (SKBs).
+Featuring diverse, natural-sounding queries that require context-specific reasoning across diverse domains, STaRK sets a new standard for assessing real-world retrieval systems.
+We contribute three large-scale retrieval datasets with human-generated queries and an automated pipeline to simulate realistic user queries.
+Our experiments on STaRK highlight significant challenges for current models in effectively handling textual and relational information.
+STaRK paves the way for future research to advance complex, multimodal retrieval systems, focusing on reducing retrieval latency and enhancing reasoning abiliites.
+Our current SKBs are limited to textual and relational information.
+Future work should incorporate additional modalities such as images, videos, and speech to provide a more comprehensive information retrieval system.
+Despite our anonymization efforts, we acknowledge that privacy remains a potential concern when extending this work to other domains with real user data, which should be protected to ensure compliance with privacy regulations.
 6 Acknowledgement
-We thank group members group members in Leskovec and Zou labs for providing valuable suggestions and conducting benchmark construction.
+We thank group members in Leskovec and Zou labs for providing valuable suggestions and conducting benchmark construction.
 We express our gratitude to the following individuals for their assistance in generating the human-generated queries (ordered by last name):
@@
 We especially thank Gavin Li at Stanford University and Zhanghan Wang at New York University for helping build the interactive interface for our SKBs.
-We also gratefully acknowledge the support of DARPA under Nos.
+We gratefully acknowledge the support of DARPA under Nos.
 N660011924033 (MCS); NSF under Nos.
 OAC-1835598 (CINES), CCF-1918940 (Expeditions), DMS-2327709 (IHBEM); Stanford Data Applications Initiative, Wu Tsai Neurosciences Institute, Stanford Institute for Human-Centered AI, Chan Zuckerberg Initiative, Amazon, Genentech, GSK, Hitachi, SAP, and UCB.
-Appendix A Mathematical Definitions of Shannon Entropy and Type-Token Ratio
-Shannon Entropy. Shannon Entropy is a measure of the uncertainty in a set of possible outcomes, quantifying the amount of information or disorder within a dataset. It is defined as follows:
-$H(X)=-\sum_{i=1}^{n}p(x_{i})\log p(x_{i})$
-where $X$ is the set of possible outcomes, $p(x_{i})$ is the probability of occurrence of the outcome $x_{i}$, and $n$ is the total number of unique outcomes.
-Higher entropy values indicate greater diversity in the distribution of outcomes.[35]
-Type-Token Ratio (TTR).
-The Type-Token Ratio is a measure of lexical diversity, calculated as the ratio of the number of unique words (types) to the total number of words (tokens) in a text.
-It is defined as follows:
-$\text{TTR}=\frac{V}{N}$
-where $V$ is the number of unique words and $N$ is the total number of words in the text. Higher TTR values indicate a higher proportion of unique words, reflecting greater lexical diversity. [41]
-Appendix B Benchmark details
+Checklist
+  1. 1.
+    For all authors…
+    1. (a)
+      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+[Yes] We clearly state our problem scope and highlight our main contribution compared to the existing works.
+    2. (b)
+      Did you describe the limitations of your work? [Yes] Please see the conclusion and future work section.
+    3. (c)
+      Did you discuss any potential negative societal impacts of your work? [Yes] Please see the conclusion and future work section.
+    4. (d)
+      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
+  2. 2.
+    If you are including theoretical results…
+    1. (a)
+      Did you state the full set of assumptions of all theoretical results? [N/A]
+    2. (b)
+      Did you include complete proofs of all theoretical results? [N/A]
+  3. 3.
+    If you ran experiments (e.g. for benchmarks)…
+    1. (a)
+      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)?
+[Yes] Our website https://stark.stanford.edu/ and GitHub codebase https://github.com/snap-stanford/STaRK contains code, data, and experimental pipelines.
+    2. (b)
+      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)?
+[Yes] Please see the experiment setup in Section 3.1, where we explained all of the choices made.
+We also make the data splits public.
+    3. (c)
+      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [N/A]
+    4. (d)
+      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)?
+[Yes] We provide the GPU device information and report the latency cost.
+  4. 4.
+    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
+    1. (a)
+      If your work uses existing assets, did you cite the creators? [Yes]
+    2. (b)
+      Did you mention the license of the assets? [Yes] We mentioned them in our webpage.
+    3. (c)
+      Did you include any new assets either in the supplemental material or as a URL? [Yes]
+    4. (d)
+      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] The resources are from exisiting public data that is open to access.
+    5. (e)
+      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content?
+[Yes] The only data with potential personally identifiable information and offensive content is Amazon semi-structured dataset, which is already made anonymized by the public resources.
+  5. 5.
+    If you used crowdsourcing or conducted research with human subjects…
+    1. (a)
+      Did you include the full text of instructions given to participants and screenshots, if applicable? [Yes]
+    2. (b)
+      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
+    3. (c)
+      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A] . We invited volunteered participants who are acknowledged.
+Appendix A Benchmark details
+A.1 Semi-structured Knowledge Bases (SKBs)
+We present the public sources that we used to construct the SKBs in the table below. We have adhered to the licenses of each public resource.
+Table 9: Sources of relational structure and textual information of the benchmarks
 relational structure textual information
@@
 gene: Ensembl, NCBI Entrez, Uniprot, UCSC, CPDB
-Table 10: Sources of relational structure and textual information of the benchmarks
-B.1 STaRK-Amazon
-Basic query templates:
+We build an interactive platform to inspect the data of all three SKBs at https://stark.stanford.edu/skb_explorer.html. We introduce more detailed data statistics below.
+Amazon SKB.
+In total, it comprises around 1.0M entities (product entities: 0.9M, brand entities: 0.1M, category entities: 1.4k, color entities: 1.7k) and 9.4M relations (also_bought: 2.8M, also_viewed: 1.9M, has_brand: 1.7M, has_category: 2.3M, has_color: 0.6M).
+MAG SKB.
+This SKB contains around 1.9M entities under four entity types (author: 1.1M, paper: 0.7M, institution: 8.7K, field_of_study: 59.5k) and 39.8M relations under four relation types (author_writes_paper: 13.5M, paper_has_field_of_study: 14.5M, paper_cites_paper: 9.7M, author_affiliated_with_institution: 2.0M).
+Prime SKB. The entity count in our knowledge base is approximately 129.3K, with around 8.1M relations. The numbers of entities in each type are listed below:
+    #disease:            17,080
+    #gene/protein:       27,671
+    #molecular_function: 11,169
+    #drug:                7,957
+    #pathway:             2,516
+    #anatomy:            14,035
+    #effect/phenotype:   15,311
+    #biological_process: 28,642
+    #cellular_component:  4,176
+    #exposure:              818
+A.2 STaRK-Amazon
+Relational query templates.
+These are the basic relational templates on STaRK-Amazon.
+Note that the final relational template can be composed of multiple basic templates.
+For example, ‘(color → product ← brand)’ represents a relational template combined from two basic relational templates.
 metapath Query template
@@
 (category → product) "What products are available in the <category> category?"
-The final relational template can be composed of multiple basic templates. For example, ‘(color → product ← brand)’ represents one such combination.
-B.2 STaRK-MAG
-Relational query templates:
+A.3 STaRK-MAG
+Relational query templates. We constructed seven relational templates below:
 metapath multi-hop query template
@@
 For multi-hop metapaths, the last metapath (institution $\rightarrow$ author $\rightarrow$ paper $\leftarrow$ field_of_study) requires an institution entity and a field_of_study entity to initialize the query.
-B.3 STaRK-Prime
-Number of entities in each type.
-#disease:            17,080
-#gene/protein:       27,671
-#molecular_function: 11,169
-#drug:                7,957
-#pathway:             2,516
-#anatomy:            14,035
-#effect/phenotype:   15,311
-#biological_process: 28,642
-#cellular_component:  4,176
-#exposure:              818
-Relational query templates.
+A.4 STaRK-Prime
+Relational query templates. For synthesized queries, we listed 28 multi-hop templates designed by experts to cover various relation types and ensure their practical relevance.
+For instance, the query “What is the drug that targets the genes or proteins expressed in <anatomy>?” serves applications in precision medicine and pharmacogenomics, aiding researchers and healthcare professionals in identifying drugs that act on genes or proteins associated with specific anatomical areas and enabling more targeted treatments.
+{mdframed}
+[backgroundcolor=gray!5, linecolor=black, linewidth=2pt, innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin=3pt, innerbottommargin=10pt]
 (effect/phenotype → [phenotype absent] → disease ← [!indication] ← drug):
@@
 (effect/phenotype → [side effect] → drug → [synergistic interaction] → drug):
-    "What drug has a synergistic interaction with the drug that has <effect/phenotype> as a side effect?",
+    "What drug has a synergistic interaction with the drug that has <effect/phenotype>
+    as a side effect?",
 (disease → [indication] → drug → [contraindication] → disease):
@@
 (biological_process → [interacts with] → gene/protein → [interacts with] → biological_process):
-    "What biological process has the common interactino pattern with gene or proteins as <biological_process>?",
+    "What biological process has the common interactino pattern with gene or proteins as
+    <biological_process>?",
 (effect/phenotype → [associated with] → gene/protein → [interacts with] → biological_process):
@@
  & (drug → [synergistic interaction] → drug):
-    "Find drugs that has a synergistic interaction with <drug> and both are indicated for the same disease.",
+    "Find drugs that has a synergistic interaction with <drug> and both are indicated
+    for the same disease.",
 (pathway → [interacts with] → gene/protein → [interacts with] → pathway)
@@
     & (gene/protein → [ppi] → gene/protein):
-    "Find gene/protein that can interect with <gene/protein> and both are associated with the same disease.",
+    "Find gene/protein that can interect with <gene/protein> and both are associated
+    with the same disease.",
 (gene/protein → [associated with] → effect/phenotype → [associated with] → gene/protein)
     & (gene/protein → [ppi] → gene/protein):
-    "Find gene/protein that can interect with <gene/protein> and both are associated with the same effect/phenotype."
+    "Find gene/protein that can interect with <gene/protein> and both are associated
+    with the same effect/phenotype."
 where []delimited-[][$\cdot$][ ⋅ ] denotes the relation type.
-Appendix C Prompts and LLM versions for Query Synthesization
+Appendix B Mathematical Definitions of Shannon Entropy and Type-Token Ratio
+Shannon Entropy. Shannon Entropy is a measure of the uncertainty in a set of possible outcomes, quantifying the amount of information or disorder within a dataset. It is defined as follows:
+$H(X)=-\sum_{i=1}^{n}p(x_{i})\log p(x_{i})$
+where $X$ is the set of possible outcomes, $p(x_{i})$ is the probability of occurrence of the outcome $x_{i}$, and $n$ is the total number of unique outcomes.
+Higher entropy values indicate greater diversity in the distribution of outcomes.[43]
+Type-Token Ratio (TTR).
+The Type-Token Ratio is a measure of lexical diversity, calculated as the ratio of the number of unique words (types) to the total number of words (tokens) in a text.
+It is defined as follows:
+$\text{TTR}=\frac{V}{N}$
+where $V$ is the number of unique words and $N$ is the total number of words in the text. Higher TTR values indicate a higher proportion of unique words, reflecting greater lexical diversity. [49]
+Appendix C Instructions for Generating Queries
+For the process of generating queries by human, the participants were given a list of entity IDs that we randomly sampled from the entire entity set.
+Then, they were asked to follow the following instructions with the support of our built interactive platform at https://stark.stanford.edu/skb_explorer.html.
+{mdframed}
+[backgroundcolor=gray!10, linecolor=black, linewidth=2pt, innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin=3pt, innerbottommargin=10pt] Task:
+1) Given the provided entity ID, review the associated document and any connected entities and multi-hop paths.
+2) Find interesting aspects of the entities by examining both their relational structures and the textual information available.
+3) Write your queries from these aspects such that the entity can satisfy all of them.
+Note:
+1) Please do not leak the name of the entity in the query.
+2) You can skip some entity IDs if you think the knowledge involved is hard to understand.
+3) Feel free to be creative with content of your queries, you can also include additional context.
+There is NO restriction on how you express the queries.
+After collecting the queries, we filtering the ground truth answers manually by human validation.
+Appendix D Experiments
+D.1 More Experimental Results
+Table 10: Positive/Non-negative rates (%) from human evaluation.
+Naturalness Diversity Practicality
+STaRK-Amazon 73.6 / 89.5 68.4 / 89.5 89.5 / 94.7
+STaRK-MAG 94.7 / 100 73.7 / 84.2 68.4 / 84.2
+STaRK-Prime 67.8 / 92.8 71.4 / 82.1 71.4 / 89.3
+Average 78.7 / 94.1 71.0 / 85.3 76.4 / 89.4
+Appendix E Prompts and LLM versions for Query Synthesization
 We summarize the LLM versions in Table 11.
-We chose these models based on their cost, how accurate they are, and whether they were the latest model during different parts of the project.
+We chose these models based on a joint consideration of their cost, how accurate they are, and whether they were the latest model during different phases of the project.
 While we used different LLMs, we checked each step separately to make sure the good quality in our benchmark datasets.
@@
 Step 4: Filtering additional answers claude-2.1, claude-2.0, claude-instant-1.2
-C.1 Extracting textual requirements
+E.1 Extracting textual requirements
 {mdframed}
@@
 Their questions tend to be straightforward, focusing on practical concerns on the symptons, effects, and inheritance etc., instead of the detailed mechanisms, which may also include more context.
-C.2 Combining relational and textual requirements
+E.2 Combining relational and textual requirements
 {mdframed}
@@
 Please output only the rewritten query:
-C.3 Filtering additional answers
+E.3 Filtering additional answers
 {mdframed}
@@
 <requirements>
-Please provide only the JSON in your response. Response:"
-Appendix D Instructions for Generating Queries
-For the process of generating queries by human, the participants were given a list of entity IDs that we randomly sampled from the entire entity set.
-Then, they were asked to follow the following instructions with the support of our built interactive demo at https://huggingface.co/spaces/snap-stanford/SKB-Explorer.
-{mdframed}[backgroundcolor=gray!10, linecolor=black, linewidth=2pt, innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin=3pt, innerbottommargin=10pt] Task:
-1) Given the provided entity ID, review the associated document and any connected entities and multi-hop paths.
-2) Find interesting aspects of the entities by examining both their relational structures and the textual information available.
-3) Write your queries from these aspects such that the entity can satisfy all of them.
-Note:
-1) Please do not leak the name of the entity in the query.
-2) You can skip some queries if you think the knowledge involved is hard to understand.
-3) Feel free to be creative with content of your queries, you can also include additional context.
-There is NO restriction on how you express the queries.
-After that, we filtering the ground truth answers manually by human validation.
+Please provide only the JSON in your response. Response: